A Complete Bibliography of Applied Statistics: 1960–1969

Nelson H. F. Beebe Center for Scientific Computing University of Utah Department of Mathematics, 322 INSCC 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                beebe@computer.org, beebe@ieee.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 4.03

Title word cross-reference

3^{2} × 2^{2} [54]. n [110]. t [92, 433, 326, 328, 385, 387, 426, 335, 295]. -Distribution [433, 326, 328, 385, 387, 335, 92]. 1 [324, 383]. 10 [348, 425, 423]. 11 [349]. 12 [350]. 13 [378]. 13th [174]. 14 [379]. 15 [380]. 16 [434, 381, 430]. 17 [382]. 18 [403]. 19 [404]. 2 [325, 402]. 20 [405]. 21 [406, 424]. 22 [417]. 23 [418]. 24 [419]. 25 [420]. 2nd [308, 357].

3 [326, 385]. 3rd [221, 358]. 4 [327, 386, 422, 426]. 4th [48]. 5 [433, 328, 387, 426]. 6 [384, 431, 429, 432, 329]. 6th [353]. 7 [384, 432, 330]. 8 [346]. 9 [347]. A. [331, 174, 333, 221, 242, 114, 231, 115, 398]. Abstracts [10, 360]. Abstracts-Statistical [10]. Acceptance [95, 63]. Access [265]. Accuracy [74]. Ackoff [397]. Actes [173]. Activity [217]. Acton [24]. Additional [228]. Additive [347]. Address [181]. Adequacy [410]. Adjustment [343]. Adler [175]. Admission [389]. Advanced [80]. Advances [316]. Advertising [224, 407]. Age [337]. Agricultural [4]. Ahamad [249]. Aid [412, 394, 375]. Aircraft [246]. Alan [80]. Algebra [277]. Algorithm [417, 431, 433, 417, 327, 326, 328, 325, 348, 385, 386, 387, 422, 425, 419, 429, 324, 403, 404, 418, 383, 330, 329, 349, 350, 421, 434, 423, 420, 347, 346, 405, 378, 379, 380, 381, 382, 430, 406, 424]. Algorithms [303, 384, 431, 433, 417, 323, 327, 326, 328, 325, 348, 385, 386, 387, 422, 425, 419, 429, 432, 324, 403, 404, 418, 383, 330, 329, 349, 350, 421, 434, 426, 423, 420, 347, 346, 405, 378, 379, 380, 428, 381, 382, 430, 406, 424]. Allocation [285, 75]. Allowing [44]. Alterations [112]. Amiel [176]. Among [53]. Anaemia [219]. Analyses [271]. Analysis [248, 297, 31, 298, 29, 168, 132, 392, 96, 204, 236, 367, 404, 299, 339, 242, 215, 54, 284, 192, 125, 6, 420, 359, 196, 253, 162, 233, 193, 380, 157, 39, 249, 209, 99, 24, 25, 305]. Analytical [87, 302]. Annotated [12]. Annual [254, 165]. Ansley [291]. Application [319, 274, 284, 196, 51, 4, 225]. Applications [57, 235, 260, 277, 339, 246, 1, 133, 39, 116, 47, 331, 90]. Applied [200, 275, 402, 333, 305, 373, 215, 139, 351, 352]. Applying [18]. Approximation [298, 205]. Arbitrary [101]. Armore [359]. Array [405]. Arrays [324]. Ascent [255]. ASCOP [264]. Aspects [2, 267, 86, 226, 59]. Assessing [162, 233]. Assessment [156, 74]. Assessors [373]. Associated [337]. Association [395]. Attitudes [220]. Audience [407]. Audley [398]. Augmented [344]. Authorship [286]. Automatic [371, 372]. Automating [169].

Automobile [53]. Auxiliary

[327, 386, 422, 426]. Available [131].

B [69, 103, 232, 280, 90, 91]. B. [399, 249]. Back [37, 70, 81, 93, 104, 117, 127, 143, 158, 178, 188, 197, 211, 314, 336, 363]. Bailey [149, 11]. Balance [372]. Balanced [31, 321]. Balancing [162, 233]. Ballade [129]. Bandwidth [193]. Barrie [278]. Bartlett [66]. Barton [102]. Based [95, 28, 75]. Basic [114]. Basis [33]. Batchelor [12]. Bayes [28]. Bayesian [237]. Behaviour [239]. Benjamin [399]. Best [204]. Between [164, 75, 219, 130]. Beyer [334]. Bhattacharya [36]. Bi [194, 393]. Bi-Spectrum [194]. Bi-Variate [393]. Bias [44, 302]. Biases [3]. Bibliography [243, 12]. Billingsley [416]. Binomial [206, 100]. Biological [277]. Biologists [209]. Biology [155, 11]. Biomathematics [232]. Birmingham [230]. Birth [247, 414]. Birth-Order [414]. Bivariate [132]. Block [31, 97]. Blocks [52]. Blood [152, 207]. Book [177, 244, 274, 155, 24, 10, 25, 35, 14, 36, 57, 47, 67, 69, 80, 102, 103, 92, 275, 331, 9, 176, 135, 136, 66, 362, 174, 333, 353, 354, 277, 221, 278, 305, 68, 156, 311, 304, 355, 308, 415, 141, 356, 358, 357, 232, 273, 306, 242, 138, 33, 290, 281, 309, 210, 113, 334, 310, 114, 137, 139, 171, 231, 416, 173, 175, 172, 280, 359, 360, 12, 351, 352, 307, 208, 48, 361, 276, 13, 79, 289, 59, 58, 157, 279, 11, 140, 154, 291, 332, 34, 115, 116, 209, 90, 91, 243]. Books [142, 177, 222, 244, 258]. Botany [1]. Britain [131, 203, 59, 279]. British [310, 7, 341]. Bronchitis [282]. Brownlee [67]. Buckland [10, 156]. Budget [22]. Buildings [202]. Bulmer [357]. Buses [407]. Business [68, 13]. Buying [293].

C [208]. C. [232, 290]. Calculation [418]. Calculations [35]. Cancer [282]. Capture [371, 373, 370, 375, 376, 372, 374]. Cards [374]. Carl [171]. Carlo [208]. Cars [203].

Case [250, 284, 216]. Cases [226, 286]. Catalytic [185]. Censored [434, 368, 381, 427, 430]. Central [328, 387, 426]. Certain [50]. Chakravarti [351, 352]. Chance [75, 102]. Characterisation [124]. Characteristic [79, 156]. Characteristics [409]. Chart [95]. Charts [33]. Check [88]. Check-List [88]. Chemical [185]. Chosen [59]. Circuit [319]. Clarifying [271]. Class [252]. Classification [28, 420, 300]. Classifying [238]. Climatological [202]. Clinical [151]. Clumping [400]. Cluster [367, 339, 380]. Coale [291]. Coarse [343]. Coarsely [100]. COBOL [141]. Cochran [353, 154]. Coefficients [170]. Collection [35]. Combinatorial [102]. Comment [247]. Comments [23, 45, 55, 77, 229, 249]. Commercial [85]. Committee [202]. Comparing [170, 5]. Comparison [88, 390, 366]. Comparisons [120]. Compensation [414]. Component [284, 365]. Components [248, 322, 133, 249]. Componets [297]. Computation [351, 393]. Computer [411, 168, 266, 412, 185, 376, 406, 424]. Computers [240, 116, 205, 115]. Computing [267, 264, 368, 427]. Concerning [206]. Conference [230, 173]. Confounding [54, 52]. Congestion [306]. Consciousness [65]. Considerations [98]. Consignments [186]. Consistency [19]. Constants [294]. Constructing [109, 321]. Construction [323, 347]. Consumer [3, 167, 220, 239, 251]. Consumers [65]. Contingency [392]. Continuous [95, 63]. Contrasts [110]. Control [364, 316, 27, 221, 33, 318, 185, 76, 48, 231]. Controlled [151]. Conventional [366]. Convergence [416]. Conversational [396]. Conway [137]. Correction [260, 233]. Correlation [170, 256]. Corrigenda [384, 385, 386, 387, 424]. Corrigendum [422, 335, 427]. Cost [377, 134, 84, 345].

Counting [345]. Counts [215, 5]. Course [276, 58, 91]. Covariance [413]. Cox [273]. Crimes [248, 296, 297, 249]. Critical [257, 50, 249]. Crop [287, 313]. Cross [84]. Cross-Section [84]. Crystals [169]. Cumulative [95, 73]. Current [27, 370]. Curve [135]. Cusum [364]. Cybernetic [210].

D [141, 273, 59, 279]. D.

[102, 103, 308, 273, 208, 332, 34]. Damped [169]. Daniel [141, 280]. Dano [57, 354]. Data [298, 254, 411, 371, 373, 434, 285, 370, 294, 375, 165, 405, 253, 100, 376, 368, 381, 427, 430, 372, 374, 24]. Data-Loggers [371]. David [102]. Deciding [286]. Decimal [257]. Decision [250, 43, 398]. Decision-Making [250]. Decisions [157, 362]. Decomposition [431, 429, 432, 329]. Defectives [95]. Deficiency [219]. Definite [432, 330]. Demand [203]. Demeny [291]. Deming [68]. Density [92]. Dependent [409]. Depot [299]. Derivation [254, 165]. Description [323]. Descriptive [137, 351]. Design [183, 31, 270, 68, 6]. Designated [112]. Designs [218, 187, 162, 233, 321]. Detecting [300]. Detection [150, 302]. Determining [169]. Developed [151]. Development [253]. Deviate [419]. Deviates [205, 36]. Diagrams [428]. Diameter [369]. Difference [164]. Differences [53, 75]. Digital [195, 116]. Dimension [324]. Direct [372]. Direct-Recording [372]. Discrete [285]. Discriminant [96]. Discrimination [285]. Discussion [268]. Disease [247]. Dispersion [301]. Disputed [286]. Distribution [35, 92, 433, 327, 326, 328, 385, 386, 387, 422, 111, 252, 50, 426, 369, 100, 335, 295, 368, 427, 153, 69, 92]. Distribution-Free [35]. Distributions [76]. Dodge [14]. Donald [69, 355]. Dose [219]. Double [14]. Dr. [149]. Draper [305].

Drawing [111]. Due [287, 313]. Duncan [221]. During [62]. Dynamic [43, 42].

E. [25, 275, 309, 307, 79, 157]. Each [97]. Earnings [86]. Easy [35]. Ecology [66]. Econometric [108, 290]. Econometrics [309]. Economic [194, 17, 120, 280, 307, 279]. Economics [9, 21, 272, 4, 281, 309, 289]. Economy [310]. Ed [221, 308]. edited [10]. Edition [174, 48, 14, 353, 358, 357, 232, 154]. Editorial [146, 180, 191, 261, 303]. Education [184, 359]. Edwards [68]. Effect [202, 218, 220, 335, 295]. Effects [418, 414, 346, 112, 162, 233]. Efficient [405]. Eilon [360]. Ekambaram [33]. Electronic [115, 205]. Elementary [391, 311, 356, 358]. Elements [401, 271]. Emanuel [47]. Empirical [272, 64]. Encoders [376]. Enders [210]. Engineering [67]. Engineers [304, 361]. England [202]. Enquiries [300]. Enquiry [61]. Enrick [48]. Epidemics [148]. Epidemiological [147]. Epidemiology [66]. Epilogue [269]. Equations [169]. Equipment [365]. Errata [207, 312, 313]. Erratum [297]. Error [168]. Errors [270, 96, 214, 121, 300]. Essentials [231]. Estimated [41]. Estimates [302, 368, 427]. Estimating [407, 161, 301, 287, 313]. Estimation [183, 108, 35, 341, 270, 434, 166, 381, 430]. Estimators [390, 322]. Evaluation [236, 403, 246]. Events [124]. Everyman [175]. Evidence [167, 64]. Evolutionary [283]. Examination [50, 59]. Example [195]. Examples [283, 271]. Exercise [299]. Expenditure [61, 214, 62]. Expenditures [341, 112]. Experiment [54]. Experimental [97, 85, 187, 164, 408]. Experiments [410, 168, 125, 110, 52, 99, 352]. Exploratory [194]. Exponential [166]. Exposure [337]. Extraneous [318].

F [14, 355, 153]. F-Distribution [153]. F. [102, 310, 90, 91]. Factor [125]. Factorial [168, 404, 54, 321, 52, 99]. Factors [125]. Factory [226]. Falconer [34]. Fallacies [27]. Family [214, 62]. Features [366]. Feinstein [176]. Feller [274]. Few [59]. Fibre [369]. Fibre-Diameter [369]. Fibreboard [226]. Field [226]. Fieldwork [415]. Figures [254, 165]. Filtering [195]. Finding [101]. Fire [202]. Fires [202]. First [415, 276, 91]. Fit [294]. Fitting [411, 394, 369, 100, 135]. Fixed [101]. Flow [72, 124]. Food [107]. Forecasting [316, 201]. Forensic [340]. Forest [371]. Forestry [375]. Foreword [234]. Forman [24]. Formation [405]. Formula [28]. Fourth [232]. Fox [307]. Fraction [377]. Fractional [110]. Frank [401, 13]. Freda [137]. Free [35]. Frequency [5]. Freund [358, 13]. Friday [401]. Front [82, 94, 105, 118, 128, 144, 159, 179, 189, 198, 212]. Function [327, 386, 422, 426, 161, 205]. Functions [92, 252, 50, 79]. Fundamentals [397]. Furneaux [59]. Further [254].

G

[14, 80, 135, 362, 353, 357, 139, 351, 352, 154]. G. [9, 278, 173, 90, 91]. Game [342, 255]. Games [331, 163]. Gamma [369]. General [267, 415]. Generating [288, 312]. Generation [205]. Genetics [34]. Geology [18]. Geophysical [196]. George [353]. Gerald [69]. Goodness [206]. Gradient [41]. Graduation [74]. Graphical [392, 394, 395]. Great [203]. Greenwood [247]. Greenwood-Yule [247]. Grohmann [171]. Group [275]. Grouped [434, 100, 368, 381, 427, 430]. Groups [8, 32, 46, 56, 78, 89, 120, 75]. Growth [161, 279]. Guest [135]. Guide [141]. Gurland [155]. Guthrie [281].

H [35, 12]. H. [305, 273]. Habits [167]. Half [166]. Hammersley [208]. Handbook [103, 334, 351, 352]. Handscomb [208]. Handwheel [19]. Hannan [275]. Harold [14, 281]. Harry [14]. Harvard [116]. Health [96, 399]. Hearing [337]. Henderson [279]. Herman [243]. Hiliary [209]. Home [44, 112]. Hooke [140]. Hope [311]. Hospital [389]. Household [61, 112]. Huitson [242, 231]. Hyper [227]. Hyper-Logarithmic [227]. Hyperbola [390]. Hypergeometric [69].

I. [351, 352]. Ideas [114]. Identification [110]. II [352, 274]. Ilersic [174]. Illustrating [255]. Improved [74]. Improvement [409]. Incentive [17]. Incidence [202]. including [277]. Incomes [341]. Incomplete [31]. Independence [392]. Index [391, 169, 229, 251]. Indices [109]. Industrial [354, 221, 360, 39, 225]. Industry [57, 98, 184, 283]. Industry-Theory [57]. Inference [80, 359, 351, 140]. Information [16, 38, 49, 71, 83, 106, 119, 145, 160, 190, 199, 213, 223, 259, 315, 388, 176, 220]. Initial [75]. Input [289]. Input-Output [289]. Inspection [292, 14, 17, 73]. Instrumental [29]. Integral [433, 326, 328, 325, 385, 387, 419, 421, 426, 393]. Integrals [327, 386, 422, 426]. Inter [62]. Inter-Week [62]. Interaction [417]. Interactions [150]. International [243, 10]. Internationale [173]. Interpolation [153]. Interpreting [107]. Interrelationships [296]. Interviewer [53, 22]. Interviewing [415, 112]. Interviews [96]. Introduction [274, 235, 260, 113, 139, 280, 359, 42, 140, 34, 57]. Inventory [30]. Inverse [205]. Inversion [432, 330]. Investigating [282, 85]. Investigation [164]. Iron [219]. Iron-Deficiency [219]. Irving [175].

J [25, 69, 275, 221, 171, 359, 13, 11, 291, 332]. J. [36, 351, 352, 307, 208, 332]. James [12]. Joan [231]. John [155, 358, 139, 360, 13].

Joint [202]. Jorgenson [332]. Journal [10]. Jr. [90, 91].

K. [67, 402, 311, 307]. Kaufmann [398]. Keen [231]. Kendall [80]. King [360]. Kingdom [182, 64]. Kitov [115]. Kreweras [173]. Krinitski [115].

L [9, 397, 306, 48, 209, 298]. L. [138]. Labour [61]. Laha [351, 352]. Language [266, 263]. Large [148, 147, 395]. Large-Scale [147]. Latin [6]. Laws [241]. Least [169]. Lector [375]. Lectures [157]. Lee [333]. Length [409]. Leontief [289]. Leser [290]. Level [99]. Lewes [310]. Lieberman [69]. Life [156, 291]. Likelihood [434, 100, 368, 381, 427, 430]. Limit [101]. Limitation [414]. Limits [126]. Line [24, 41, 265, 185]. Linear [57, 410, 4]. Lines [394]. Linguistics [50]. Linkage [111, 367, 380]. List [88, 131]. Local [8, 32, 46, 56, 78, 89]. Location [299]. Log [369]. Log-Normal [369]. Logarithmic [227]. Logarithms [257]. Loggers [371]. Logistic [161]. London [407]. Loss [337, 287, 313]. Lot [63]. Lot-Size [63]. Loveday [276, 58]. Lowell [361]. Lower [101]. LSD [293]. Lukacs [79]. Lung [282].

M [36, 333, 308, 310, 351, 352, 208]. M. [35, 176, 66, 397, 357]. M1 [21]. Macdonald [304]. Machine [19]. Magic [162, 233]. Magnetic [376]. Magnetic-Tape [376]. Main [346]. Maintenance [220]. Making [250, 398]. Malinvaud [309]. Manager [362]. Managerial [362, 309]. Manufacturers [217]. Marginal [403, 84]. Market [235, 260, 238]. Marketing [241, 236, 237, 250]. Marvin [136]. Mathematical [240, 309, 216, 130]. Mathematics [304]. Matrices [256]. Matrix [431, 277, 429, 432, 330, 329, 317, 349, 350]. Matter [37, 70, 81, 93, 104, 82, 94, 105, 117, 127, 143, 118, 128, 144, 158, 178, 188, 159, 179, 189, 197, 211, 198, 212, 314, 336, 363]. Maurice [80]. Maximum [434, 134, 100, 368, 381, 427, 430]. Maximum-Likelihood [368, 427]. McCall [332]. McCracken [141]. McIntosh [308]. Meaning [293]. Means [403, 420]. Measurement [224]. Measurements [40, 343, 369]. Measures [416]. Measuring [217]. Mechanism [409]. Media [240]. Medicine [155]. Meetings [8, 32, 46, 56, 78, 89]. Mensch [331]. Metallurgical [133]. Method [248, 297, 247, 44, 344, 109, 74, 339, 166, 301, 249, 99, 300, 10]. Methodology [67]. Methods [108, 35, 254, 30, 135, 353, 278, 415, 281, 309, 351, 352, 161, 1, 321, 287, 313, 11, 345, 208, 355, 415]. METO [266]. Military [157]. Miller [273]. Mills [382]. Minimum [367, 378, 379]. Ministry [202, 61]. Miscellanea [296, 298, 254, 342, 344, 413, 396, 323, 168, 255, 322, 412, 252, 343, 299, 228, 250, 394, 257, 377, 206, 170, 414, 218, 256, 220, 169, 253, 271, 288, 301, 219, 287, 395, 229, 302, 251, 272, 249, 345, 205, 300]. Missing [31, 6]. Mite [287, 313]. Mixed [107, 166, 99]. Mode [396]. Model [50, 291]. Models [148, 108, 155, 147, 410, 240, 282, 66, 411, 354, 239]. Modern [47, 358, 13]. Moments [166]. Monde [172]. Monte [208]. Moore [362]. Morbidity [109]. Morlat [173]. Morrison [355]. Mortality [282, 120]. Moses [138]. Mosteller [90, 91]. Motor [64]. Mrs [296, 309]. Multi [270, 346, 365]. Multi-Component [365]. Multi-Stage [270]. Multi-Way [346]. Multidimensional [324]. Multiple [129, 131, 204, 317, 253, 408]. Multivariate [355, 236, 320, 1, 39, 130, 209]. Multiway [294].

N [102]. N. [92, 305, 115]. NBD [293]. NBD/LSD [293]. Negative [322, 100].

News [245]. Nikhilesh [36]. Nine [52].

Noise [337]. Non

[410, 328, 387, 322, 426, 377, 335, 295, 51]. Non-Central [328, 387, 426]. Non-Linear [410]. Non-Negative [322]. Non-Normality [335, 295]. Non-Parametric [51]. Non-Sampling [377]. Noncentral [433]. Norbert [48]. Normal [36, 325, 419, 111, 320, 421, 434, 369, 301, 393, 122, 368, 381, 427, 430, 205]. Normality [335, 295]. Normalizing [349]. Norman [11]. Not-at-Home [44]. Note [296, 303, 7, 342, 377, 149, 288, 312]. Notes [23, 45, 55, 77, 41]. Notices [245]. Number [95, 251]. Numbers [391]. Numerical [135, 374].

O [243]. Observational [413].

Observations [31, 6]. Observed [5]. Occurrence [75]. Office [202]. Old [151]. On-Line [265, 185]. One [168, 324, 321]. One-Way [321]. Open [88]. Opening [181]. Operation [3, 283]. Operational [29, 366, 409]. Opérationelle [172, 173]. Operations [397, 12]. Optimal [183, 377, 332]. Optimum [101, 340, 64]. Oral [219]. Order [247, 414, 218, 162, 233]. Ore [200]. Organization [202]. Orthogonal [348, 425]. Output [289]. Owen [69, 103]. Owner [112].

P [305]. P. [135, 362, 304, 113, 416, 279]. Package [383]. Panel [3]. Paper [283, 149]. Parameter [108, 390]. Parameters [368, 427]. Parametric [51]. Particular [21, 340]. Party [415]. Parzen [47]. Patrol [292]. Paul [291]. Peatman [139]. Percentages [121]. Permutations [288, 312, 138]. Pharmaceutical [98]. Phenomena [147]. Physical [30]. Piece [241]. Pinsker [176]. Plan [63]. Planning [410, 240, 352]. Pleasure [9]. Plots [406, 424]. Plotting [320]. Poisson [164]. Policy [307, 332]. Polynomials [348, 425].

Population [66, 182, 335, 295]. Populations [148, 122, 291]. Port [374]. Port-a-Punch [374]. Portable [376]. Portfolio [216]. Positive [432, 330]. Postal [86, 300]. Potential [3]. Practical [293, 226]. Practice [171]. Precision [41]. Prediction [206, 218]. Preference [2, 107]. Preliminary [392]. Presence [247, 414]. Price [65, 220, 251]. Prices [229]. Principal [248, 297, 284, 133, 249]. Principles [357, 171]. Printing [379]. Probabilities [225]. Probability [274, 47, 69, 362, 334, 416, 175, 401, 90, 91, 275]. Problem [200, 152, 207, 124, 340, 251]. Problèmes [172]. Problems [27, 107, 228, 18, 123, 290]. Procedure [95, 264]. Procedures [28, 370]. Proceedings [306, 116]. Process [185]. Processes [85, 243, 176, 273]. Processing [375]. Product [250]. Production [354, 63]. Products [350]. Professional [86]. Profitability [345]. Program [396, 294]. Programme [168, 125]. Programmes [131]. Programming [261, 262, 200, 57, 267, 265, 264, 266, 141, 263, 43, 269, 268, 216, 42, 4, 130, 113]. Project [246, 243]. Projects [98]. Prologue [262]. Properties [218]. Proportional [225]. Provided [17]. Psychology [359]. Publication [15, 26]. Publications [60]. Punch [374]. Purchasers [53]. Purchasing [200, 239].

Quade [157]. Qualities [107]. Quality [364, 221, 74, 33, 231, 76, 48]. Quantal [294]. Quantitative [280, 294, 307, 34]. Quarterly [254, 165]. Quasi [392]. Quasi-Independence [392]. Quenouille [35]. Questioning [88]. Queue [409]. Queue-Length-Dependent [409]. Queueing [333]. Queues [366, 409]. Quick [255].

R [10, 174, 277, 156, 273, 360]. R.

[397, 351, 352, 361, 140, 398, 90, 91]. R1 [383]. R10 [428]. R12 [429]. R2 [421]. R21 [430]. R27 [431]. R3 [423]. R4 [425]. R44 [432]. R56 [433]. R70 [434]. R8 [426]. Radner [332]. RAND [157]. Random [36, 176, 138, 210, 400, 288, 312, 75, 205]. Rapid [35]. Ratio [382]. Reactor [319]. Received [15, 26, 60, 142, 177, 222, 244, 258]. Recherche [173, 172]. Reciprocal [382]. Recorded [62]. Recording [372, 374]. Records [151]. Red [287, 313]. Reduce [151]. Refereeing [342]. Reference [21, 340]. Reflections [182]. Refractive [169]. Regional [291]. Regions [183]. Regression [25, 129, 132, 305, 131, 204, 299, 317, 170, 87, 253, 271]. Regulating [366]. Related [2]. Relation [101]. Relations [130]. Relationship [219, 80]. Relevance [147]. Reliability [319, 136]. Remark [431, 433, 425, 429, 432, 383, 421, 434, 426, 423, 430, 431, 433, 425, 429, 432, 383, 421, 434, 426, 423, 428, 430]. Repairs [112]. Repeat [293]. Repeat-Buying [293]. Replacement [332, 365]. Replicates [110]. Reply [296]. Report [415]. Reporting [96, 272]. Representation [275]. Reproductive [414]. Resale [220]. Research [235, 260, 98, 371, 236, 237, 202, 338, 12, 397, 68, 280]. Respondents [238, 112, 300]. Response [88, 219]. Responses [107, 408]. Restricted [405]. Restrictions [228]. Retail [229]. Review [240, 230, 370]. Reviews [177, 244, 274, 155, 24, 10, 25, 35, 14, 36, 57, 47, 67, 69, 80, 102, 103, 92, 275, 331, 9, 176, 135, 136, 66, 362, 397, 174, 402, 333, 353, 354, 277, 221, 278, 305, 68, 156, 311, 304, 355, 308, 415, 141, 356, 358, 357, 232, 273, 306, 242, 138, 33, 399, 290, 281, 309, 210, 113, 334, 310, 114, 137, 139, 171, 231, 416, 173, 175, 172, 280, 359, 401, 400, 360, 12, 351, 352, 307, 208, 48, 361, 276, 13, 79, 289, 59, 58, 157, 279, 11, 140, 154, 291, 332, 34, 115, 116, 398].

Reviews [209, 90, 91, 243]. Risk [282]. Roach [400]. Roads [21]. Robert [276, 58]. Robinson [210]. Role [193]. Romig [14]. Rourke [90, 91]. Routine [383, 253]. Roy [351, 352]. Royal [8, 32, 46, 56, 78, 89]. Rules [365].

S [24, 9, 176, 66, 157, 34]. S. [277, 33, 359, 400]. Safety [319]. Sales [201]. Sample [44, 68, 301]. Sampled [335, 295]. Samples [395]. Sampling [95, 14, 30, 344, 270, 182, 72, 356, 17, 377, 73, 214, 186, 63, 154, 122, 225, 114, 401]. Samuel [360]. Sasieni [397]. Scale [147, 428, 406, 424]. Scatter [428]. Scheduling [360]. Schemes [73]. Science [67, 340, 398]. Sciences [277]. Scientific [114, 140]. Scientists [402, 304, 361]. Scoring [151]. Seal [209]. Searle [277]. Second [14, 218, 58, 154]. Section [84]. Sections [8, 32, 46, 56, 78, 89]. Selection [98, 59, 51, 406, 424]. Selections [216, 428]. Self [366]. Self-Regulating [366]. Semi [432, 330, 372]. Semi-Automatic [372]. Semi-Definite [432, 330]. Sengupta [36, 307]. Separable [200]. Sequential [278, 122]. Series [194, 215, 243]. Server [409]. Service [409]. Set [204]. Setting [19]. Several [125]. Severe [219]. Shackle [9]. Short [131, 201]. Short-Term [201]. Significance [108, 164]. Silvey [309]. Simon [50]. Simple [301, 395]. Simplex [283]. Simulating [324]. Simulation [124]. Simultaneous [108]. Single [14, 367, 409, 380, 365]. Single-Server [409]. Singular [317]. Size [151, 63, 225]. Small [148, 217]. Smirnov [92]. Smith [305, 232, 306]. Snedecor [353]. Social [402]. Society [8, 32, 46, 56, 78, 89]. Socio [120]. Socio-Economic [120]. Soil [338]. Solve [169]. Some [148, 41, 2, 316, 267, 98, 27, 3, 339, 202, 167, 370, 86, 283, 18, 123, 226, 1, 162, 233, 59, 229, 133, 39, 249]. Space [150]. Space-Time [150]. Spanning [367, 378, 379]. Specified [134]. Spectral [192, 196, 193]. Spectrum [194]. Spider [287, 313]. Spread [202]. Square [6]. Squares [350, 169, 162, 233]. Stability [176]. Stable [291]. Stage [270]. Standard [121]. Standardization [318]. Standards [40]. State [182]. Statistical [8, 32, 46, 56, 78, 89, 261, 262, 303, 384, 431, 10, 35, 67, 103, 2, 136, 126, 267, 433, 417, 265, 264, 385, 386, 387, 422, 425, 235, 260, 266, 419, 163, 98, 353, 156, 355, 429, 432, 263, 403, 404, 418, 383, 184, 33, 421, 434, 426, 281, 309, 423, 420, 359, 123, 351, 269, 405, 268, 378, 379, 380, 428, 11, 302, 381, 382, 430, 406, 424, 345, 209, 90]. Statistics [7, 80, 389, 362, 29, 174, 265, 402, 277, 221, 311, 304, 308, 412, 72, 358, 357, 343, 237, 20, 246, 373, 338, 18, 175, 280, 351, 352, 340, 361, 276, 58, 51, 130, 310, 278, 399, 334, 137, 139, 171, 13]. Steepest [255]. Stochastic [148, 155, 147, 66, 273, 243]. Stock [27]. Storage [405]. Straight [24, 41, 394]. Straight-Line [24]. Strategy [85]. Stuart [80, 114]. Student [326, 385]. Studies [413, 309, 284, 75, 408]. Study [3, 194, 250, 84, 216, 354]. Sub [204, 383]. Sub-Routine [383]. Sub-Set [204]. Substantial [75]. Suggested [109]. Suitable [205]. Suits [280]. Sum [95]. Sums [350, 73]. Survey [192, 238, 22, 86, 214, 112, 62]. Surveys [44, 270, 352, 225]. Sven [57, 354]. Symbolic [113]. Symmetric [431, 429, 432, 330, 329, 349]. Symposium [306, 116]. System [151, 411, 366]. Systems [88, 389, 265, 157, 210].

T [11]. Table [404, 257, 347, 346]. Tables [14, 36, 69, 92, 392, 138, 334, 294, 291, 153, 103]. Tail [241]. Tail-Piece [241]. Tape [376]. Target [101, 76]. Taro [356]. Taste [20]. Tastes [167]. Taxes [64]. Tea [287, 313]. Teacher [308]. Teaching [412]. Team [255, 243]. Technical [86].

Technique [318]. Techniques [364, 331, 255, 290, 351, 154]. Technology [202]. Term [201]. Terms [168]. Test [410, 340, 395]. Testing [152, 207, 20, 414, 97, 256, 123, 35]. Tests [2]. Their [27, 116]. Theoretical [354]. Theory [319, 274, 10, 57, 47, 67, 80, 331, 136, 293, 273, 43, 400, 307, 225, 333, 356, 306]. There [97]. Thomas [90, 91]. Thorbecke [307]. Three [97, 287, 313]. Tiers [172]. Time [194, 215, 150, 243]. Tolerance [126]. Tool [19, 87]. Top [255]. Tops [369]. Towns [341]. Trace [373]. Traffic [72]. Transformation [218]. Transit [186]. Transportation [228]. Treatment [391, 27]. Treatments [168, 97]. Tree [378, 379]. Trees [367]. Trend [321]. Trends [227, 337]. Trial [151]. Trials [226]. Triangular [431, 429, 432, 329, 405]. Troisieme [173]. Truncated [122, 51]. Truncation [218]. Two [390, 392, 257, 284, 97, 125, 5, 164, 62]. Two-Decimal [257]. Two-Parameter [390]. Two-Way [392]. Two-Week [62].

Uncertainty [250]. Unemployment [7]. United [182, 64]. Units [97]. University [59]. Upon [95]. Upper [101]. Use [240, 282, 362, 413, 348, 425, 371, 415, 72, 299, 409, 373, 73, 271, 162, 233]. Usefulness [293, 187]. Useless [132]. Using [30, 151, 125].

V [92, 290]. Validation [107]. Value [101]. Variability [22, 61]. Variable [63]. Variables [176, 206, 318, 164]. Variance [31, 168, 322, 404, 125, 218, 6, 420, 162, 233, 242]. Variate [393]. Variations [202, 62]. Visual [394]. Vital [399]. Vol [351, 352]. Volume [16, 38, 49, 71, 83, 106, 119, 145, 160, 190, 199, 213, 223, 259, 315, 388, 274, 80, 402, 232, 309].

W [397, 353, 281, 332]. W.

[10, 68, 156, 306, 334, 59, 154]. Wales [202]. Walker [296]. Wassily [289]. Wavelets [210]. Way [392, 346, 321]. Ways [255]. Week [62]. Wegner [113]. Wetherill [278]. Which [123]. Wildfowl [215]. Wilkinson [306]. William [274, 353]. Williams [25, 13]. Wine [361]. Wold [243]. Wool [369]. Work [272]. Workbook [311]. Working [415].

Yamane [356]. Yeomans [402]. Yield [134]. Yule [247].

Zelen [136].

References

Pearce:1960:SAM

[1] S. Clifford Pearce and Derek A. Holland. Some applications of multivariate methods in botany. Applied Statistics, 9(1): 1–7, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bliss:1960:SSA

[2] C. I. Bliss. Some statistical aspects of preference and related tests. Applied Statistics, 9(1):8–19, March 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1960:SSP

[3] A. S. C. Ehrenberg. A study of some potential biases in the operation of a consumer panel. Applied Statistics, 9(1): 20–27, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vincent:1960:ALP

[4] Shirely E. Vincent. An application of linear programming to agricultural economics. Applied Statistics, 9(1):28– 36, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kupperman:1960:CTO

[5] Morton Kupperman. On comparing two observed frequency counts. Applied Statistics, 9(1):37–42, March 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kramer:1960:AVL

[6] Clyde Young Kramer and Suzanne Glass. Analysis of variance of a latin square design with missing observations. Applied Statistics, 9(1):43–50, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1960:NBU

[7] Donald Bailey. Note on British unemployment statistics. Applied Statistics, 9(1):51–59, March 1960. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:MSLa

[8] Anonymous. Meetings of sections and local groups of the royal statistical society. Applied Statistics, 9(1):60–61, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Booker:1960:BRE

[9] H. S. Booker. Book reviews: Economics for Pleasure, by G. L. S. Shackle. Applied Statistics, 9(1):62–63, March 1960.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1960:BRI

[10] D. G. Beech. Book reviews: International Journal of Abstracts-Statistical Theory and Method, edited by W. R. Buckland. Applied Statistics, 9(1):63– 64, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1960:BRS

[11] Cedric A. B. Smith. Book reviews: Statistical Methods in Biology, by Norman T. J. Bailey. Applied Statistics, 9(1): 64–66, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muir:1960:BRO

[12] A. Muir. Book reviews: Operations Research: An Annotated Bibliography, by James H. Batchelor. Applied Statistics, 9 (1):66, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ridley:1960:BRM

[13] T. M. Ridley. Book reviews: Modern Business Statistics, by John E. Freund and Frank J. Williams. Applied Statistics, 9(1):67, March 1960. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1960:BRS

[14] D. G. Beech. Book reviews: Sampling Inspection Tables, Single and Double Sampling. Second Edition, by Harold F. Dodge and Harry G. Romig. Applied Statistics, 9(1):67–68, March 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:PRa

[15] Anonymous. Publication received. Applied Statistics, 9(1):68, March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:VIPa

[16] Anonymous. Volume information. Applied Statistics, 9(1):[4], March 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1960:EIP

[17] I. D. Hill. The economic incentive provided by sampling inspection. Applied Statistics, 9(2):69–81, June 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krumbein:1960:SPA

[18] W. C. Krumbein. Some problems in applying statistics to geology. Applied Statistics, 9(2):82–91, June 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1960:CSM

[19] E. Nigel Corlett and Geoffrey Gregory. The consistency of setting of a machine tool handwheel. Applied Statistics, 9(2): 92–102, June 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gridgeman:1960:STT

[20] Norman T. Gridgeman. Statistics and taste testing. Applied Statistics, 9(2): 103–112, June 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burdess:1960:ENR

[21] T. K. Burdess. The economics of new roads with particular reference to M1. Applied Statistics, 9(2):113–121, June 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1960:IVB

[22] W. F. F. Kemsley. Interviewer variability and a budget survey. Applied Statistics, 9(2):122–128, June 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:NC

[23] Anonymous. Notes and comments. Applied Statistics, 9(2):129, June 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1960:BRA

[24] D. G. Beech. Book reviews: Analysis of Straight-Line Data, by Forman S. Acton. Applied Statistics, 9(2):130– 132, June 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1960:BRRa

[25] D. G. Beech. Book reviews: Regression Analysis, by E. J. Williams. Applied Statistics, 9(2):130–132, June 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:PRb

[26] Anonymous. Publication received. Applied Statistics, 9(2):132, June 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Duckworth:1960:SCP

[27] Eric Duckworth. Stock control problems: Some fallacies in their current treatment. Applied Statistics, 9(3):133– 151, November 1960. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Birnbaum:1960:CPB

[28] A. Birnbaum and A. E. Maxwell. Classification procedures based on Bayes’s formula. Applied Statistics, 9(3):152–169, November 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Calder:1960:OSI

[29] Allan B. Calder. Operational statistics in instrumental analysis. Applied Statistics, 9(3):170–177, November 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bryson:1960:PIU

[30] Marion R. Bryson. Physical inventory using sampling methods. Applied Statistics, 9(3):178–188, November 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baird:1960:AVB

[31] Hugh Robert Baird and Clyde Young Kramer. Analysis of variance of a balanced incomplete block design with missing observations. Applied Statistics, 9(3):189–198, November 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:MSLb

[32] Anonymous. Meetings of sections and local groups of the royal statistical society. Applied Statistics, 9(3):199–200, November 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1960:BRS

[33] I. D. Hill. Book reviews: The Statistical Basis of Quality Control Charts, by S. K. Ekambaram. Applied Statistics, 9(3): 201–202, November 1960. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Veale:1960:BRI

[34] A. M. O. Veale. Book reviews: Introduction to Quantitative Genetics, by D. S. Falconer. Applied Statistics, 9(3): 202–203, November 1960. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1960:BRRb

[35] D. G. Beech. Book reviews: Rapid Statistical Calculations: A Collection of Distribution-Free and Easy Methods of Estimation and Testing, by M. H. Quenouille. Applied Statistics, 9(3):203–204, November 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1960:BRT

[36] D. G. Beech. Book reviews: Tables of Random Normal Deviates, by J. M. Sengupta and Nikhilesh Bhattacharya. Applied Statistics, 9(3):204, November 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:BMP

[37] Anonymous. Back matter. Applied Statistics, 9(3):[1], November 1960. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1960:VIPb

[38] Anonymous. Volume information. Applied Statistics, 9(3):[1], November 1960. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:1961:SIA

[39] B. A. M. Thomas. Some industrial applications of multivariate analysis. Applied Statistics, 10(1):1–8, March 1961. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conway:1961:MS

[40] Freda Conway. Measurements and standards. Applied Statistics, 10(1):9– 13, March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1961:SNP

[41] Donald G. Beech. Some notes on the precision of the gradient of an estimated straight line. Applied Statistics, 10(1): 14–31, March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simpson:1961:IDP

[42] M. G. Simpson. An introduction to dynamic programming. Applied Statistics, 10(1):32–38, March 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1961:DPD

[43] D. V. Lindley. Dynamic programming and decision theory. Applied Statistics, 10(1):39–51, March 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartholomew:1961:MAH

[44] D. J. Bartholomew. A method of allowing for ‘not-at-home’ bias in sample surveys. Applied Statistics, 10(1): 52–59, March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:NCa

[45] Anonymous. Notes and comments. Applied Statistics, 10(1):60, March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:MSLa

[46] Anonymous. Meetings of sections and local groups of the royal statistical society. Applied Statistics, 10(1):61–62, March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1961:BRM

[47] D. G. Beech. Book reviews: Modern Probability Theory and Its Applications, by Emanuel Parzen. Applied Statistics, 10(1):63–64, March 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pridmore:1961:BRQ

[48] W. A. Pridmore. Book reviews: Quality Control. (4th Edition), by Norbert L. Enrick. Applied Statistics, 10(1): 64, March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:VIPa

[49] Anonymous. Volume information. Applied Statistics, 10(1):[4], March 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Herdan:1961:CES

[50] Gustav Herdan. A critical examination of Simon’s model of certain distribution functions in linguistics. Applied Statistics, 10(2):65–76, June 1961. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tallis:1961:ANP

[51] G. M. Tallis. An application of nonparametric statistics to truncated selection. Applied Statistics, 10(2):77– 82, June 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkie:1961:CFE

[52] D. Wilkie. Confounding in 3 3 factorial experiments in nine blocks. Applied Statistics, 10(2):83–92, June 1961. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jung:1961:IDA

[53] Allen F. Jung. Interviewer differences among automobile purchasers. Applied Statistics, 10(2):93–97, June 1961. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Inkson:1961:AFE

[54] R. H. E. Inkson. The analysis of a 3^{2}×2^{2} factorial experiment with confounding. Applied Statistics, 10(2):98–107, June 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:NCb

[55] Anonymous. Notes and comments. Applied Statistics, 10(2):108–109, June 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:MSLb

[56] Anonymous. Meetings of sections and local groups of the royal statistical society. Applied Statistics, 10(2):110–111, June 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1961:BRL

[57] D. G. Beech. Book reviews: Linear Programming in Industry-Theory and Applications: An Introduction, by Sven Dano. Applied Statistics, 10(2): 112, June 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandon:1961:BRS

[58] Frank Sandon. Book reviews: A second course in statistics, by Robert Loveday. Applied Statistics, 10(2):112–113, June 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandon:1961:BRC

[59] Frank Sandon. Book reviews: The Chosen Few: An Examination of Some Aspects of University Selection in Britain, by W. D. Furneaux. Applied Statistics, 10(2):113–116, June 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:PR

[60] Anonymous. Publications received. Applied Statistics, 10(2):116, June 1961.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1961:HEE

[61] W. F. F. Kemsley. The household expenditure enquiry of the Ministry of Labour: Variability in the 1953–54 enquiry. Applied Statistics, 10(3):117–135, November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Turner:1961:IWV

[62] Robert Turner. Inter-week variations in expenditure recorded during a two-week survey of family expenditure. Applied Statistics, 10(3):136–146, November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Read:1961:VLS

[63] D. R. Read and D. W. Beattie. The variable lot-size acceptance sampling plan for continuous production. Applied Statistics, 10(3):147–156, November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walters:1961:EEO

[64] A. A. Walters. Empirical evidence on optimum motor taxes for the United Kingdom. Applied Statistics, 10(3):157–169, November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gabor:1961:PCC

[65] Andre Gabor and C. W. J. Granger. On the price consciousness of consumers. Applied Statistics, 10(3):170– 188, November 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bulmer:1961:BRS

[66] M. G. Bulmer. Book reviews: Stochastic Population Models in Ecology and Epidemiology, by M. S. Bartlett. Applied Statistics, 10(3):189, November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1961:BRS

[67] D. G. Beech. Book reviews: Statistical Theory and Methodology in Science and Engineering, by K. A. Brownlee. Applied Statistics, 10(3):189–190, November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Emmett:1961:BRS

[68] B. P. Emmett. Book reviews: Sample Design in Business Research, by W. Edwards Deming. Applied Statistics, 10(3): 190–191, November 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1961:BRT

[69] D. G. Beech. Book reviews: Tables of the Hypergeometric Probability Distribution, by Gerald J. Lieberman and Donald B. Owen. Applied Statistics, 10(3): 191–192, November 1961. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:BMP

[70] Anonymous. Back matter. Applied Statistics, 10(3):[1], November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1961:VIPb

[71] Anonymous. Volume information. Applied Statistics, 10(3):[1], November 1961. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garwood:1962:SUT

[72] F. Garwood. The sampling and use of traffic flow statistics. Applied Statistics, 11(1):1–15, March 1962. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1962:UCS

[73] Kenneth W. Kemp. The use of cumulative sums for sampling inspection schemes. Applied Statistics, 11(1):16– 31, March 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fuller:1962:QAG

[74] A. G. Fuller and H. H. Collins. Quality assessment: A graduation method for improved accuracy. Applied Statistics, 11(1):32–46, March 1962. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Radhakrishna:1962:COS

[75] S. Radhakrishna and Ian Sutherland. The chance occurrence of substantial initial differences between groups in studies based on random allocation. Applied Statistics, 11(1):47–54, March 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morrison:1962:TDQ

[76] S. J. Morrison. Target distributions for quality control. Applied Statistics, 11(1): 55–63, March 1962. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:NC

[77] Anonymous. Notes and comments. Applied Statistics, 11(1):64, March 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:MSLa

[78] Anonymous. Meetings of sections and local groups of the royal statistical society. Applied Statistics, 11(1):65–66, March 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosenhead:1962:BRC

[79] J. V. Rosenhead. Book reviews: Characteristic Functions, by E. Lukacs. Applied Statistics, 11(1):67, March 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1962:BRA

[80] D. G. Beech. Book reviews: The Advanced Theory of Statistics. Volume 2: Inference and Relationship, by Maurice G. Kendall and Alan Stuart. Applied Statistics, 11(1):67–68, March 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:BMPa

[81] Anonymous. Back matter. Applied Statistics, 11(1):[6], March 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:FMPa

[82] Anonymous. Front matter. Applied Statistics, 11(1):[2], March 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:VIPa

[83] Anonymous. Volume information. Applied Statistics, 11(1):[4], March 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oliver:1962:CSS

[84] F. R. Oliver. A cross-section study of marginal cost. Applied Statistics, 11(2): 69–78, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Klingel:1962:ESI

[85] A. R. Klingel, Jr. and R. G. McIntyre. An experimental strategy for investigating commercial processes. Applied Statistics, 11(2):79–92, June 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1962:STA

[86] W. F. F. Kemsley. Some technical aspects of a postal survey into professional earnings. Applied Statistics, 11(2):93– 105, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moore:1962:RAT

[87] P. G. Moore. Regression as an analytical tool. Applied Statistics, 11(2): 106–119, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Belson:1962:CCL

[88] William Belson and Judith A. Duncan. A comparison of the check-list and the open response questioning systems. Applied Statistics, 11(2):120–132, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:MSLb

[89] Anonymous. Meetings of sections and local groups of the royal statistical society. Applied Statistics, 11(2):133–134, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wishart:1962:BRPa

[90] D. M. G. Wishart. Book reviews: Probability With Statistical Applications, by F. Mosteller, R. E. K. Rourke and G. B. Thomas Jr. Applied Statistics, 11(2): 135–136, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wishart:1962:BRPb

[91] D. M. G. Wishart. Book reviews: Probability; a First Course, by F. Mosteller, R. E. K. Rourke and G. B. Thomas Jr. Applied Statistics, 11(2):135–136, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1962:BRT

[92] D. G. Beech. Book reviews: Tables for the Distribution and Density Functions of t-Distribution, by N. V. Smirnov. Applied Statistics, 11(2):136, June 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:BMPb

[93] Anonymous. Back matter. Applied Statistics, 11(2):[6], June 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:FMPb

[94] Anonymous. Front matter. Applied Statistics, 11(2):[2], June 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beattie:1962:CAS

[95] D. W. Beattie. A continuous acceptance sampling procedure based upon a cumulative sum chart for the number of defectives. Applied Statistics, 11(3): 137–147, November 1962. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fisher:1962:DAR

[96] Gordon Fisher. A discriminant analysis of reporting errors in health interviews. Applied Statistics, 11(3):148–163, November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1962:TTT

[97] Peter W. M. John. Testing two treatments when there are three experimental units in each block. Applied Statistics, 11(3):164–169, November 1962. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1962:SSC

[98] O. L. Davies. Some statistical considerations in the selection of research projects in the pharmaceutical industry. Applied Statistics, 11(3):170–183, November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkie:1962:MAM

[99] D. Wilkie. A method of analysis of mixed level factorial experiments. Applied Statistics, 11(3):184– 195, November 1962. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OCarroll:1962:FNB

[100] F. M. O’Carroll. Fitting a negative binomial distribution to coarsely grouped data by maximum likelihood. Applied Statistics, 11(3):196–201, November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bettes:1962:FOT

[101] D. C. Bettes. Finding an optimum target value in relation to a fixed lower limit and an arbitrary upper limit. Applied Statistics, 11(3):202–210, November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1962:BRC

[102] D. G. Beech. Book reviews: Combinatorial Chance, by F. N. David and D. E. Barton. Applied Statistics, 11 (3):211, November 1962. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1962:BRH

[103] D. G. Beech. Book reviews: Handbook of Statistical Tables, by D. B. Owen. Applied Statistics, 11(3):211–212, November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:BMPc

[104] Anonymous. Back matter. Applied Statistics, 11(3):[8], November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:FMPc

[105] Anonymous. Front matter. Applied Statistics, 11(3):[2], November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1962:VIPb

[106] Anonymous. Volume information. Applied Statistics, 11(3):[1], November 1962. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gregson:1963:VPI

[107] R. A. M. Gregson. Validation problems in interpreting preference responses to mixed food qualities. Applied Statistics, 12(1):1–13, March 1963. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ball:1963:SSM

[108] R. J. Ball. The significance of simultaneous methods of parameter estimation in econometric models. Applied Statistics, 12(1):14–25, March 1963. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Draper:1963:SMC

[109] G. J. Draper. A suggested method for constructing indices of morbidity. Applied Statistics, 12(1):26–37, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1963:ICF

[110] J. A. Nelder. Identification of contrasts in fractional replicates of 2 n experiments. Applied Statistics, 12(1):38– 43, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Edwards:1963:LDN

[111] A. W. F. Edwards. A linkage for drawing the normal distribution. Applied Statistics, 12(1):44–45, March 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neter:1963:EID

[112] John Neter and Joseph Waksberg. Effects of interviewing designated respondents in a household survey of home owners’ expenditures on alterations and repairs. Applied Statistics, 12(1):46– 60, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1963:BRI

[113] J. Aneurin John. Book reviews: An Introduction to Symbolic Programming, by P. Wegner. Applied Statistics, 12(1): 61–62, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1963:BRB

[114] G. G. W. Kalton. Book reviews: Basic Ideas of Scientific Sampling, by A. Stuart. Applied Statistics, 12(1):62, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wegner:1963:BRE

[115] Peter Wegner. Book reviews: Electronic Computers, by A. I. Kitov and N. A. Krinitski. Applied Statistics, 12 (1):63, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wegner:1963:BRP

[116] Peter Wegner. Book reviews: Proceedings of a Harvard Symposium on Digital Computers and their Applications. Applied Statistics, 12(1):63–64, March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:BMPa

[117] Anonymous. Back matter. Applied Statistics, 12(1):[6], March 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:FMPa

[118] Anonymous. Front matter. Applied Statistics, 12(1):[2], March 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:VIPa

[119] Anonymous. Volume information. Applied Statistics, 12(1):[3], March 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kilpatrick:1963:MCS

[120] S. J. Kilpatrick. Mortality comparisons in socio-economic groups. Applied Statistics, 12(2):65–86, June 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stuart:1963:SEP

[121] Alan Stuart. Standard errors for percentages. Applied Statistics, 12(2):87– 101, June 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swamy:1963:SST

[122] P. S. Swamy. On sequential sampling from truncated normal populations. Applied Statistics, 12(2):102– 106, June 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mitchell:1963:TWS

[123] Jeremy Mitchell. Testing for which?: Some statistical problems. Applied Statistics, 12(2):107–112, June 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Knowles:1963:CFE

[124] E. A. G. Knowles and D. S. Stewart. Characterisation of the flow of events — A problem of simulation. Applied Statistics, 12(2):113–128, June 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1963:UTF

[125] Peter W. M. John. Using a two factor analysis of variance programme for experiments with several factors. Applied Statistics, 12(2):129–132, June 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burrows:1963:STL

[126] G. L. Burrows. Statistical tolerance limits — what are they? Applied Statistics, 12(2):133–144, June 1963. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:BMPb

[127] Anonymous. Back matter. Applied Statistics, 12(2):[6], June 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:FMPb

[128] Anonymous. Front matter. Applied Statistics, 12(2):[2], June 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1963:BMR

[129] Tom Corlett. Ballade of multiple regression. Applied Statistics, 12(3):145, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wegner:1963:RBM

[130] P. Wegner. Relations between multivariate statistics and mathematical programming. Applied Statistics, 12(3): 146–150, November 1963. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garside:1963:SLM

[131] M. J. Garside. A short list of multiple regression programmes available in Britain. Applied Statistics, 12(3):151– 160, November 1963. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1963:BRA

[132] A. S. C. Ehrenberg. Bivariate regression analysis is useless. Applied Statistics, 12(3):161–179, November 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Spurrell:1963:SMA

[133] D. J. Spurrell. Some metallurgical applications of principal components. Applied Statistics, 12(3):180–188, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Michaels:1963:MYS

[134] S. E. Michaels and P. J. Pengilly. Maximum yield for specified cost. Applied Statistics, 12(3):189–193, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Buckland:1963:BRN

[135] W. R. Buckland. Book reviews: Numerical Methods of Curve Fitting, by P. G. Guest. Applied Statistics, 12(3):194, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Buckland:1963:BRS

[136] W. R. Buckland. Book reviews: Statistical Theory of Reliability, by Marvin Zelen. Applied Statistics, 12(3):194–195, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1963:BRD

[137] G. Kalton. Book reviews: Descriptive Statistics, by Freda Conway. Applied Statistics, 12(3):195–196, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1963:BRT

[138] M. J. R. Healy. Book reviews: Tables of Random Permutations, by L. E. Moses. Applied Statistics, 12(3):196, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1963:BRI

[139] G. Kalton. Book reviews: Introduction to Applied Statistics, by John G. Peatman. Applied Statistics, 12(3):197, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1963:BRI

[140] T. M. F. Smith. Book reviews: Introduction to Scientific Inference, by R. Hooke. Applied Statistics, 12(3):198, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garside:1963:BRG

[141] M. J. Garside. Book reviews: A guide to COBOL programming, by Daniel D. McCracken. Applied Statistics, 12(3): 198–199, November 1963. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:BR

[142] Anonymous. Books received. Applied Statistics, 12(3):200, November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:BMPc

[143] Anonymous. Back matter. Applied Statistics, 12(3):[7], November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:FMPc

[144] Anonymous. Front matter. Applied Statistics, 12(3):[2], November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1963:VIPb

[145] Anonymous. Volume information. Applied Statistics, 12(3):[2], November 1963. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:Ea

[146] Anonymous. Editorial. Applied Statistics, 13(1):1, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartlett:1964:RSM

[147] M. S. Bartlett. The relevance of stochastic models for large-scale epidemiological phenomena. Applied Statistics, 13(1):2– 8, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1964:SSM

[148] Norman T. J. Bailey. Some stochastic models for small epidemics in large populations. Applied Statistics, 13(1):9–19, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1964:NDB

[149] R. W. Morgan. A note on Dr. Bailey’s paper. Applied Statistics, 13(1):20–24, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Knox:1964:DST

[150] E. G. Knox and M. S. Bartlett. The detection of space-time interactions. Applied Statistics, 13(1):25–30, 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carpenter:1964:USS

[151] R. G. Carpenter, Ann Petrie, and E. Rosemary Dalton. Using a scoring system developed from old records to reduce the size of a controlled clinical trial. Applied Statistics, 13(1):31–42, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Finucan:1964:BTP

[152] H. M. Finucan. The blood testing problem. Applied Statistics, 13(1):43–50, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See errata [207].

Zinger:1964:ITF

[153] A. Zinger. On interpolation in tables of the F-distribution. Applied Statistics, 13(1):51–53, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1964:BRS

[154] T. M. F. Smith. Book reviews: Sampling Techniques, Second Edition, by W. G. Cochran. Applied Statistics, 13(1):54, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartlett:1964:BRS

[155] M. S. Bartlett. Book reviews: Stochastic Models in Medicine and Biology, by John Gurland. Applied Statistics, 13 (1):54–55, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Epstein:1964:BRS

[156] Benjamin Epstein. Book reviews: Statistical Assessment of The Life Characteristic, by W. R. Buckland. Applied Statistics, 13(1):56, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Selwidge:1964:BRR

[157] J. Selwidge. Book reviews: RAND lectures on systems analysis: Analysis for military decisions, by E. S. Quade. Applied Statistics, 13(1):56, 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:BMPa

[158] Anonymous. Back matter. Applied Statistics, 13(1):[6], 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:FMPa

[159] Anonymous. Front matter. Applied Statistics, 13(1):[4], 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:VIPa

[160] Anonymous. Volume information. Applied Statistics, 13(1):[4], 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oliver:1964:MEL

[161] F. R. Oliver. Methods of estimating the logistic growth function. Applied Statistics, 13(2):57–66, 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Phillips:1964:UMS

[162] J. P. N. Phillips. The use of magic squares for balancing and assessing order effects in some analysis of variance designs. Applied Statistics, 13(2):67–73, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Daly:1964:SG

[163] C. Daly. Statistical games. Applied Statistics, 13(2):74–83, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ractliffe:1964:SDB

[164] J. F. Ractliffe. The significance of the difference between two Poisson variables: An experimental investigation. Applied Statistics, 13(2):84–86, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lisman:1964:DQF

[165] J. H. C. Lisman and J. Sandee. Derivation of quarterly figures from annual data. Applied Statistics, 13(2):87–90, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Joffe:1964:MEE

[166] A. D. Joffe. Mixed exponential estimation by the method of half moments. Applied Statistics, 13(2):91–98, 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:1964:SEC

[167] David S. Huang. Some evidence on consumer tastes and habits. Applied Statistics, 13(2):99–109, 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coulter:1964:MCP

[168] B. S. Coulter. Miscellanea: A computer programme for analysis of variance in experiments with factorial treatments and one or more error terms. Applied Statistics, 13(2):110–117, 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Matz:1964:MAD

[169] A. W. Matz. Miscellanea: Automating damped least squares to solve the equations determining refractive index of crystals. Applied Statistics, 13(2): 118–127, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

James:1964:MCR

[170] J. W. James. Miscellanea: Comparing regression and correlation coefficients. Applied Statistics, 13(2):127–132, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1964:BRP

[171] G. Kalton. Book reviews: Principles and Practice of Statistics, by Carl J. Grohmann. Applied Statistics, 13 (2):133, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:1964:BRR

[172] J. A. Lambert. Book reviews: Recherche Opérationelle et Problèmes du Tiers Monde. Applied Statistics, 13(2):133, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:1964:BRA

[173] J. A. Lambert. Book reviews: Actes de la Troisieme Conference Internationale de Recherche Operationelle, by G. Kreweras and G. Morlat. Applied Statistics, 13(2):133–134, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1964:BRS

[174] C. Chatfield. Book reviews: Statistics.

(13th Edition), by A. R. Ilersic. Applied Statistics, 13(2):134, 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:1964:BRP

[175] J. A. Lambert. Book reviews: Probability and Statistics for Everyman, by Irving Adler. Applied Statistics, 13(2):134, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brillinger:1964:BRI

[176] David R. Brillinger. Book reviews: Information and Information Stability of Random Variables and Processes, by M. S. Pinsker and Amiel Feinstein. Applied Statistics, 13(2):134–135, 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:BRB

[177] Anonymous. Book reviews: Books received. Applied Statistics, 13(2): 135–136, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:BMPb

[178] Anonymous. Back matter. Applied Statistics, 13(2):[5], 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:FMPb

[179] Anonymous. Front matter. Applied Statistics, 13(2):[2], 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1964:Eb

[180] Anonymous. Editorial. Applied Statistics, 13(3):137, 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Cox:1964:OA

[181] D. R. Cox and W. J. Jennett. Opening address. Applied Statistics, 13(3): 138–145, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Emmett:1964:RSP

[182] B. P. Emmett. Reflections on the state of population sampling in the United Kingdom. Applied Statistics, 13 (3):146–157, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Aitchison:1964:EOD

[183] J. Aitchison. The estimation of optimal design regions. Applied Statistics, 13(3):158–173, 1964. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Hall:1964:SEI

[184] J. A. P. Hall and W. A. Pridmore. Statistical education for industry. Applied Statistics, 13(3):174–194, 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Lee:1964:LCC

[185] W. T. Lee. On-line computer control of a catalytic chemical process. Applied Statistics, 13(3):195–207, 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Page:1964:STC

[186] D. J. Page, O. Feldman, and D. E. Edwards. The sampling of transit consignments. Applied Statistics, 13(3): 208–220, 1964. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Michaels:1964:UED

[187] S. E. Michaels. The usefulness of experimental designs. Applied Statistics, 13 (3):221–235, 1964. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Anonymous:1964:BMPc

[188] Anonymous. Back matter. Applied Statistics, 13(3):[3], 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Anonymous:1964:FMPc

[189] Anonymous. Front matter. Applied Statistics, 13(3):[2], 1964. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Anonymous:1964:VIPb

[190] Anonymous. Volume information. Applied Statistics, 13(3):[2], 1964. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Proceedings of the September 1963 Joint Conference of the Research and Industrial Applications Sections of the Royal Statistical Society.

Anonymous:1965:E

[191] Anonymous. Editorial. Applied Statistics, 14(1):1, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jenkins:1965:SSA

[192] G. M. Jenkins. A survey of spectral analysis. Applied Statistics, 14(1):2–32, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Priestley:1965:RBS

[193] M. B. Priestley. The role of bandwidth in spectral analysis. Applied Statistics, 14(1):33–47, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Godfrey:1965:ESB

[194] M. D. Godfrey. An exploratory study of the bi-spectrum of economic time series. Applied Statistics, 14(1):48–69, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Alavi:1965:EDF

[195] A. S. Alavi and G. M. Jenkins. An example of digital filtering. Applied Statistics, 14(1):70–74, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neidell:1965:GAS

[196] N. S. Neidell. A geophysical application of spectral analysis. Applied Statistics, 14(1):75–88, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1965:BMPa

[197] Anonymous. Back matter. Applied Statistics, 14(1):[6], 1965. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1965:FMPa

[198] Anonymous. Front matter. Applied Statistics, 14(1):[2], 1965. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1965:VIPa

[199] Anonymous. Volume information. Applied Statistics, 14(1):[5], 1965. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beale:1965:SPA

[200] E. M. L. Beale, P. J. Coen, and A. D. J. Flowerdew. Separable programming applied to an ore purchasing problem. Applied Statistics, 14(2/3):89–101, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harrison:1965:STS

[201] P. J. Harrison. Short-term sales forecasting. Applied Statistics, 14(2/3):102–139, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hogg:1965:MTF

[202] Jane M. Hogg. Ministry of Technology and Fire Offices’ Committee Joint Fire Research Organization: The effect of some climatological variations on the incidence and spread of fires in buildings in England and Wales from 1951 to 1961. Applied Statistics, 14 (2/3):140–161, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OHerlihy:1965:DCG

[203] C. St. J. O’Herlihy. Demand for cars in Great Britain. Applied Statistics, 14 (2/3):162–195, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garside:1965:BSS

[204] M. J. Garside. The best sub-set in multiple regression analysis. Applied Statistics, 14(2/3):196–200, 1965. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wetherill:1965:MAI

[205] G. B. Wetherill. Miscellanea: An approximation to the inverse normal function suitable for the generation of random normal deviates on electronic computers. Applied Statistics, 14(2/3): 201–205, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:1965:MGP

[206] David S. Huang. Miscellanea: Goodness of prediction concerning binomial variables. Applied Statistics, 14(2/3): 206–209, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Finucan:1965:EBT

[207] H. M. Finucan. Errata: The blood testing problem. Applied Statistics, 14 (2/3):210, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [152].

Prescott:1965:BRM

[208] P. Prescott. Book reviews: Monte Carlo Methods, by J. M. Hammersley and D. C. Handscomb. Applied Statistics, 14 (2/3):211, 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wetherill:1965:BRM

[209] G. B. Wetherill. Book reviews: Multivariate Statistical Analysis for Biologists, by Hiliary L. Seal. Applied Statistics, 14(2/3):211–212, 1965. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jenkins:1965:BRR

[210] G. M. Jenkins. Book reviews: Random Wavelets and Cybernetic Systems, by Enders A. Robinson. Applied Statistics, 14(2/3):212–213, 1965. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1965:BMPb

[211] Anonymous. Back matter. Applied Statistics, 14(2/3):[5], 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1965:FMPb

[212] Anonymous. Front matter. Applied Statistics, 14(2/3):[2], 1965. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1965:VIPb

[213] Anonymous. Volume information. Applied Statistics, 14(2/3):[1], 1965. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1966:SEF

[214] W. F. F. Kemsley. Sampling errors in the family expenditure survey. Applied Statistics, 15(1):1–14, 1966. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holgate:1966:TSA

[215] P. Holgate. Time series analysis applied to wildfowl counts. Applied Statistics, 15(1):15–23, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paine:1966:CSM

[216] Neil R. Paine. A case study in mathematical programming of portfolio selections. Applied Statistics, 15(1):24–36, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hanna:1966:MAS

[217] Frank A. Hanna. Measuring the activity of small manufacturers. Applied Statistics, 15(1):37–47, 1966. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnson:1966:MPS

[218] A. F. Johnson. Miscellanea: Properties of second order designs: Effect of transformation or truncation on prediction variance. Applied Statistics, 15(1):48– 50, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rowland:1966:MRB

[219] H. A. K. Rowland. Miscellanea: The relationship between the dose of oral iron and the response in severe irondeficiency anaemia. Applied Statistics, 15(1):50–56, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Langaskens:1966:MEI

[220] Y. Langaskens and W. van Rijckeghem.

Miscellanea: The effect of information on consumers’ attitudes toward resale price maintenance. Applied Statistics, 15(1):56–62, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ellis:1966:BRQ

[221] D. M. Ellis. Book reviews: Quality Control and Industrial Statistics. 3rd Ed, by A. J. Duncan. Applied Statistics, 15(1):63–64, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1966:BR

[222] Anonymous. Books received. Applied Statistics, 15(1):64, 1966. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1966:VIP

[223] Anonymous. Volume information. Applied Statistics, 15(1):[4], 1966. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Benjamin:1966:MA

[224] B. Benjamin. The measurement of advertising. Applied Statistics, 15(2):65– 73, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vanBeeck:1966:ATS

[225] J. G. van Beeck and J. B. Vermetten. Application of the theory of sampling with probabilities proportional to size on industrial surveys. Applied Statistics, 15(2):74–93, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:1966:SPA

[226] D. J. Page. Some practical aspects of factory and field trials with fibreboard cases. Applied Statistics, 15(2):94–109, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fenske:1966:HLT

[227] R. W. Fenske. Hyper-logarithmic trends.

Applied Statistics, 15(2):110–115, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Haley:1966:MTP

[228] K. B. Haley and A. J. Smith. Miscellanea: Transportation problems with additional restrictions. Applied Statistics, 15(2):116–127, 1966. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1966:MSC

[229] T. M. F. Smith. Miscellanea: Some comments on the index of retail prices. Applied Statistics, 15(2):128–135, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Edwards:1966:BCR

[230] D. E. Edwards. 1965 birmingham conference review. Applied Statistics, 15 (2):136–148, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1966:BRE

[231] K. W. Kemp. Book reviews: Essentials of Quality Control, by A. Huitson and Joan Keen. Applied Statistics, 15(2):149, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffiths:1966:BRB

[232] J. D. Griffiths. Book reviews: Biomathematics, Volume 1. Fourth Edition, by C. A. B. Smith. Applied Statistics, 15 (2):149–150, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Phillips:1966:CUM

[233] J. P. N. Phillips. Correction: The use of magic squares for balancing and assessing order effects in some analysis of variance designs. Applied Statistics, 15(2):151, 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

T:1966:F

[234] T. C. and A. S. C. E. Foreword. Applied Statistics, 15(3):153–154, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Corlett:1966:SAM

[235] T. Corlett. Statistical applications in market research: An introduction. Applied Statistics, 15(3):155–156, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Gatty:1966:MAM

[236] Ronald Gatty. Multivariate analysis for marketing research: An evaluation. Applied Statistics, 15(3):157–172, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Green:1966:BSM

[237] Paul E. Green and Ronald E. Frank. Bayesian statistics and marketing research. Applied Statistics, 15(3):173– 190, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Joyce:1966:CMS

[238] T. Joyce and C. Channon. Classifying market survey respondents. Applied Statistics, 15(3):191–215, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Lawrence:1966:MCP

[239] Raymond J. Lawrence. Models of consumer purchasing behaviour. Applied Statistics, 15(3):216–233, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Broadbent:1966:MPC

[240] Simon Broadbent. Media planning and computers by 1970: A review of the use of mathematical models in media planning. Applied Statistics, 15(3):234– 256, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Ehrenberg:1966:LMT

[241] A. S. C. Ehrenberg. Laws in marketing: A tail-piece. Applied Statistics, 15 (3):257–267, March 1966. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Hay:1966:BRA

[242] W. A. Hay. Book reviews: The Analysis of Variance, by A. Huitson. Applied Statistics, 15(3):268, March 1966. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Wishart:1966:BRB

[243] D. M. G. Wishart. Book reviews: Bibliography on Time Series and Stochastic Processes: An International Team Project, by Herman O. A. Wold. Applied Statistics, 15(3):269–270, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Anonymous:1966:BRB

[244] Anonymous. Book reviews: Books received. Applied Statistics, 15(3): 270, March 1966. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Anonymous:1966:NN

[245] Anonymous. News and notices. Applied Statistics, 15(3):271, March 1966. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). Research into Marketing.

Harvey:1967:ASA

[246] R. A. Harvey. Applications of statistics to aircraft project evaluation. Applied Statistics, 16(1):1–12, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barker:1967:BOP

[247] D. J. P. Barker and R. G. Record. Birth order and the presence of disease: A comment on the Greenwood-Yule method. Applied Statistics, 16 (1):13–16, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ahamad:1967:ACM

[248] B. Ahamad. An analysis of crimes by the method of principal components. Applied Statistics, 16(1):17–35, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See comments [249, 297, 296].

Walker:1967:MSC

[249] Monica A. Walker. Miscellanea: Some critical comments on “An Analysis of Crimes by the Method of Principal Components” by B. Ahamad. Applied Statistics, 16(1):36–39, 1967. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [248, 297, 296].

Harris:1967:MNP

[250] Lawrence Harris. Miscellanea: New product marketing: A case study in decision-making under uncertainty. Applied Statistics, 16(1):39–42, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thakurta:1967:MPC

[251] B. K. Guha Thakurta. Miscellanea:

On a problem in consumers’ price index number. Applied Statistics, 16(1):42–45, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gittus:1967:MCD

[252] J. H. Gittus. Miscellanea: On a class of distribution functions. Applied Statistics, 16(1):45–50, 1967. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Newton:1967:MDM

[253] R. G. Newton and D. J. Spurrell. Miscellanea: A development of multiple regression for the analysis of routine data. Applied Statistics, 16(1):51–64, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Boot:1967:MFM

[254] J. C. G. Boot, W. Feibes, and J. H. C.

Lisman. Miscellanea: Further methods of derivation of quarterly figures from annual data. Applied Statistics, 16(1):65–75, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farlie:1967:MQW

[255] D. J. Farlie and Joan Keen. Miscellanea: Quick ways to the top: A team game illustrating steepest ascent techniques. Applied Statistics, 16(1):75–80, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kullback:1967:MTC

[256] S. Kullback. Miscellanea: On testing correlation matrices. Applied Statistics, 16(1):80–85, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1967:MCT

[257] M. J. R. Healy. Miscellanea: A critical table of two-decimal logarithms. Applied Statistics, 16(1):86, 1967. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1967:BRP

[258] Anonymous. Books received. Applied Statistics, 16(1):[1], 1967. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1967:VIP

[259] Anonymous. Volume information. Applied Statistics, 16(1):[3], 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1967:CSA

[260] T. Corlett. Correction: Statistical applications in market research: An introduction. Applied Statistics, 16(1):[1], 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1967:SPE

[261] Anonymous. Statistical programming:

Editorial. Applied Statistics, 16(2):87, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1967:SPP

[262] Anonymous. Statistical programming:

Prologue. Applied Statistics, 16(2):88, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1967:SPS

[263] J. C. Gower, H. R. Simpson, and A. H.

Martin. Statistical programming: A statistical programming language. Applied Statistics, 16(2):89–99, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:1967:SPA

[264] B. E. Cooper. Statistical programming: ASCOP — A statistical computing procedure. Applied Statistics, 16 (2):100–110, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Colin:1967:SPL

[265] A. J. T. Colin. Statistical programming: On-line access systems in statistics. Applied Statistics, 16(2):111–119, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Craddock:1967:SPM

[266] J. M. Craddock and M. H. Freeman. Statistical programming: The METO computer language. Applied Statistics, 16 (2):120–123, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chambers:1967:SPS

[267] John M. Chambers. Statistical programming: Some general aspects of statistical computing. Applied Statistics, 16 (2):124–132, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:1967:SPD

[268] E. S. Page, M. J. R. Healy, J. N. R. Jeffers, A. Baines, and G. M. Paddle. Statistical programming: Discussion. Applied Statistics, 16(2):133–148, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1967:SPE

[269] J. A. Nelder and B. E. Cooper. Statistical programming: Epilogue. Applied Statistics, 16(2):149–151, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Durbin:1967:DMS

[270] J. Durbin. Design of multi-stage surveys for the estimation of sampling errors. Applied Statistics, 16(2):152–164, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Newton:1967:MEU

[271] R. G. Newton and D. J. Spurrell. Miscellanea: Examples of the use of elements for clarifying regression analyses. Applied Statistics, 16(2):165–172, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:1967:MRE

[272] J. J. Thomas. Miscellanea: The reporting of empirical work in economics. Applied Statistics, 16(2):172–176, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffiths:1967:BRT

[273] J. D. Griffiths. Book reviews: The Theory of Stochastic Processes, by D. R. Cox and H. D. Miller. Applied Statistics, 16(2):177, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartholomew:1967:BRI

[274] D. J. Bartholomew. Book reviews: An Introduction to Probability Theory and Its Application. Volume II, by William Feller. Applied Statistics, 16(2):177–179, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berry:1967:BRG

[275] G. Berry. Book reviews: Group Representation and Applied Probability, by E. J. Hannan. Applied Statistics, 16(2):179, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ractliffe:1967:BRF

[276] J. F. Ractliffe. Book reviews: A first course in statistics, by Robert Loveday. Applied Statistics, 16(2):179–180, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Draper:1967:BRM

[277] G. J. Draper. Book reviews: Matrix Algebra for the Biological Sciences (including Applications in Statistics), by S. R. Searle. Applied Statistics, 16(2):180, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ellis:1967:BRS

[278] D. M. Ellis. Book reviews: Sequential Methods in Statistics, by G. Barrie Wetherill. Applied Statistics, 16 (2):180–181, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shone:1967:BRE

[279] Robert Shone. Book reviews: Economic Growth in Britain, by P. D. Henderson. Applied Statistics, 16(2):181–182, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lewis:1967:BRS

[280] J. Parry Lewis. Book reviews: Statistics: An Introduction to Quantitative Economic Research, by Daniel B. Suits. Applied Statistics, 16(2):183, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoyle:1967:BRS

[281] M. H. Hoyle. Book reviews: Statistical Methods in Economics, by Harold W. Guthrie. Applied Statistics, 16 (2):183–184, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Buck:1967:MUI

[282] S. F. Buck and A. J. Wicken. Models for use in investigating the risk of mortality from lung cancer and bronchitis. Applied Statistics, 16(3):185–210, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kenworthy:1967:SES

[283] I. C. Kenworthy. Some examples of simplex evolutionary operation in the paper industry. Applied Statistics, 16 (3):211–224, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1967:TCS

[284] J. N. R. Jeffers. Two case studies in the application of principal component analysis. Applied Statistics, 16(3):225–236, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hills:1967:DAD

[285] M. Hills. Discrimination and allocation with discrete data. Applied Statistics, 16 (3):237–250, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sarndal:1967:DCD

[286] Carl-Erik Särndal. On deciding cases of disputed authorship. Applied Statistics, 16(3):251–268, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sen:1967:MTM

[287] A. R. Sen and R. P. Chakrabarty. Miscellanea: Three methods of estimating the loss of tea crop due to red spider mite. Applied Statistics, 16(3):269–272, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See errata [313].

Page:1967:MNG

[288] E. S. Page. Miscellanea: A note on generating random permutations. Applied Statistics, 16(3):273–274, 1967. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See errata [312].

Round:1967:BRI

[289] J. I. Round. Book reviews: Input-Output Economics, by Wassily Leontief. Applied Statistics, 16(3):275, 1967. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoyle:1967:BRE

[290] M. H. Hoyle. Book reviews: Econometric Techniques and Problems, by C. E. V. Leser. Applied Statistics, 16(3):276, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thompson:1967:BRR

[291] Eric J. Thompson. Book reviews: Regional Model Life Tables and Stable Populations, by Ansley J. Coale and Paul Demeny. Applied Statistics, 16 (3):276–277, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beattie:1968:PI

[292] D. W. Beattie. Patrol inspection. Applied Statistics, 17(1):1–16, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1968:PMU

[293] A. S. C. Ehrenberg. The practical meaning and usefulness of the NBD/LSD theory of repeat-buying. Applied Statistics, 17(1):17–32, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lewis:1968:PFC

[294] J. A. Lewis. A program to fit constants to multiway tables of quantitative and quantal data. Applied Statistics, 17(1):33–41, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ratcliffe:1968:EDN

[295] J. F. Ratcliffe. The effect on the t distribution of non-normality in the sampled population. Applied Statistics, 17(1):42– 48, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [335].

Ahamad:1968:MNI

[296] B. Ahamad. Miscellanea: A note on the interrelationships of crimes — A reply to Mrs Walker. Applied Statistics, 17(1):49–51, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [248, 249, 297].

Anonymous:1968:EAC

[297] Anonymous. Erratum: An analysis of crimes by the method of principal componets. Applied Statistics, 17(1):51, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [248, 249, 296].

Barrodale:1968:MAA

[298] I. Barrodale. Miscellanea: L 1 approximation and the analysis of data. Applied Statistics, 17(1):51–57, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffiths:1968:MUR

[299] D. O. Griffiths. Miscellanea: The use of regression analysis in a depot location exercise. Applied Statistics, 17(1):57–63, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

leRoux:1968:MMD

[300] A. A. le Roux. Miscellanea: A method of detecting errors of classification by respondents to postal enquiries. Applied Statistics, 17(1):64–69, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1968:MSM

[301] P. Prescott. Miscellanea: A simple method of estimating dispersion from normal sample. Applied Statistics, 17 (1):70–74, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stark:1968:MSD

[302] A. E. Stark. Miscellanea: Statistical detection of bias in analytical estimates. Applied Statistics, 17(1):74–78, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1968:SAE

[303] Anonymous. Statistical algorithms: Editorial note. Applied Statistics, 17(1):79– 82, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1968:BRMa

[304] B. E. Evans. Book reviews: Mathematics and Statistics for Scientists and Engineers, by P. Macdonald. Applied Statistics, 17(1):83, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ellis:1968:BRA

[305] D. M. Ellis. Book reviews: Applied Regression Analysis, by N. P. Draper and H. Smith. Applied Statistics, 17(1):83– 84, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffiths:1968:BRP

[306] J. D. Griffiths. Book reviews: Proceedings of a Symposium on Congestion Theory, by W. L. Smith and W. E. Wilkinson. Applied Statistics, 17(1):84, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oliver:1968:BRT

[307] F. R. Oliver. Book reviews: The Theory of Quantitative Economic Policy, by K. A. Fox, J. K. Sengupta and E. Thorbecke. Applied Statistics, 17(1):85–86, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1968:BRS

[308] I. G. Evans. Book reviews: Statistics for the Teacher, 2nd ed, by D. M. McIntosh. Applied Statistics, 17(1):86, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoyle:1968:BRS

[309] M. H. Hoyle. Book reviews: Statistical Methods of Econometrics. Volume 6 of Studies in Mathematical and Managerial Economics, by E. Malinvaud and Mrs A. Silvey. Applied Statistics, 17(1):87, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1968:BRS

[310] D. Adams Jones. Book reviews: Statistics of the British Economy, by F. M. M.

Lewes. Applied Statistics, 17(1):87–88, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1968:BRE

[311] I. G. Evans. Book reviews: Elementary Statistics, a Workbook, by K. Hope. Applied Statistics, 17(1):88, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:1968:ENG

[312] E. S. Page. Errata: A note on generating random permutations. Applied Statistics, 17(1):89, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [288].

Sen:1968:ETM

[313] A. R. Sen and R. P. Chakrabarty. Errata: Three methods of estimating the loss of tea crop due to red spider mite. Applied Statistics, 17(1):89, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [287].

Anonymous:1968:BMPa

[314] Anonymous. Back matter. Applied Statistics, 17(1):[1], 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1968:VIP

[315] Anonymous. Volume information. Applied Statistics, 17(1):[3], 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Box:1968:SRA

[316] G. E. P. Box and G. M. Jenkins. Some recent advances in forecasting and control. Applied Statistics, 17(2):91–109, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1968:MRS

[317] M. J. R. Healy. Multiple regression with a singular matrix. Applied Statistics, 17 (2):110–117, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1968:STC

[318] G. Kalton. Standardization: A technique to control for extraneous variables. Applied Statistics, 17(2):118–136, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Babik:1968:ART

[319] S. Babik. Application of reliability theory to a reactor safety circuit. Applied Statistics, 17(2):137–156, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1968:MNP

[320] M. J. R. Healy. Multivariate normal plotting. Applied Statistics, 17 (2):157–161, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Phillips:1968:MCO

[321] J. P. N. Phillips. Methods of constructing one-way and factorial designs balanced for trend. Applied Statistics, 17 (2):162–170, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Federer:1968:MNN

[322] W. T. Federer. Miscellanea: Nonnegative estimators for components of variance. Applied Statistics, 17(2): 171–174, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Computing:1968:MCD

[323] The Working Party on Statistical Computing. Miscellanea: The construction and description of algorithms. Applied Statistics, 17(2):175–179, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1968:AAS

[324] J. C. Gower. Algorithms: Algorithm AS 1: Simulating multidimensional arrays in one dimension. Applied Statistics, 17 (2):180–185, June 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/1. See remark

[383].

Cooper:1968:AAN

[325] B. E. Cooper. Algorithms: Algorithm AS 2: The normal integral. Applied Statistics, 17(2):186– 188, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/2. See remark [421].

Cooper:1968:AAIa

[326] B. E. Cooper. Algorithms: Algorithm AS 3: The integral of Student’s tdistribution. Applied Statistics, 17(2): 189–190, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/3.

Cooper:1968:AAA

[327] B. E. Cooper. Algorithms: Algorithm AS 4: An auxiliary function for distribution integrals. Applied Statistics, 17(2): 190–192, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/4. See remarks [426].

Cooper:1968:AAIb

[328] B. E. Cooper. Algorithms: Algorithm AS 5: The integral of the non-central t-distribution. Applied Statistics, 17 (2):193–194, June 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/5. See remarks

[327, 426, 433].

Healy:1968:AAT

[329] M. J. R. Healy. Algorithms: Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 17 (2):195–197, June 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/6. See corrigen-

dum [384] and remark [429, 432, 431].

Healy:1968:AAI

[330] M. J. R. Healy. Algorithms: Algorithm AS 7: Inversion of a positive semi-definite symmetric matrix. Applied Statistics, 17(2):198–199, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/7.

See corrigendum [384] and remark [432].

Blofeld:1968:BRT

[331] R. E. Blofeld. Book reviews: Theory of Games: Techniques and Applications, by A. Mensch. Applied Statistics, 17(2):200, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomson:1968:BRO

[332] F. C. Thomson. Book reviews: Optimal Replacement Policy, by D. W. Jorgenson, J. J. McCall and D. Radner. Applied Statistics, 17(2):200, 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1968:BRA

[333] A. Davies. Book reviews: Applied Queueing Theory, by A. M. Lee. Applied Statistics, 17(2):201, 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1968:BRH

[334] D. Adams Jones. Book reviews: Handbook of Tables for Probability and Statistics, by W. E. Beyer. Applied Statistics, 17(2):201–202, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ratcliffe:1968:CED

[335] J. F. Ratcliffe. Corrigendum: The effect on the t-distribution of non-normality in the sampled population. Applied Statistics, 17(2):203, 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [295].

Anonymous:1968:BMPb

[336] Anonymous. Back matter. Applied Statistics, 17(2):[1], 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pearson:1968:THL

[337] J. C. G. Pearson and P. Sprent. Trends in hearing loss associated with age or exposure to noise. Applied Statistics, 17 (3):205–215, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Inkson:1968:SSR

[338] R. H. E. Inkson. Statistics in soil research. Applied Statistics, 17(3): 216–225, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harrison:1968:MCA

[339] P. J. Harrison. A method of cluster analysis and some applications. Applied Statistics, 17(3):226–236, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Parker:1968:OTS

[340] J. B. Parker and A. Holford. Optimum test statistics with particular reference to a forensic science problem. Applied Statistics, 17(3):237–251, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1968:EIE

[341] W. E. Cox, Jr. The estimation of incomes and expenditures in British towns. Applied Statistics, 17(3):252–259, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chambers:1968:MNG

[342] J. M. Chambers and Agnes M. Herzberg.

Miscellanea: A note on the game of refereeing. Applied Statistics, 17(3): 260–263, 1968. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:1968:MAS

[343] P. L. Goldsmith. Miscellanea: Adjustment of statistics from coarse measurements. Applied Statistics, 17(3): 264–270, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatterjee:1968:MMA

[344] S. Chatterjee. Miscellanea: The method of augmented sampling. Applied Statistics, 17(3):271–274, 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1968:MCC

[345] D. H. Ward. Miscellanea: “Counting the Cost” — statistical methods and profitability. Applied Statistics, 17 (3):274–276, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1968:AAM

[346] J. A. Nelder. Algorithms: Algorithm AS 8: Main effects from a multi-way table. Applied Statistics, 17(3):277–279, September 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/8.

Nelder:1968:AAC

[347] J. A. Nelder. Algorithms: Algorithm AS 9: Construction of additive table. Applied Statistics, 17(3):279–283, September 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/9.

Cooper:1968:AAU

[348] B. E. Cooper. Algorithms: Algorithm AS 10: The use of orthogonal polynomials. Applied Statistics, 17(3):283–287, September 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/10. See remarks [423, 425].

Herraman:1968:AAN

[349] Carol Herraman. Algorithms: Algorithm AS 11: Normalizing a symmetric matrix. Applied Statistics, 17(3): 287–288, September 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/11.

Herraman:1968:AAS

[350] Carol Herraman. Algorithms: Algorithm AS 12: Sums of squares and products matrix. Applied Statistics, 17(3): 289–292, September 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/12.

Murphy:1968:BRHa

[351] B. P. Murphy. Book reviews: Handbook of Methods of Applied Statistics. Vol. I: Techniques of Computation, Descriptive Methods and Statistical Inference, by I. M. Chakravarti, R. G. Laha and J. Roy. Applied Statistics, 17(3):293–294, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Murphy:1968:BRHb

[352] B. P. Murphy. Book reviews: Handbook of Methods of Applied Statistics. Vol. II: Planning of Surveys and Experiments, by I. M. Chakravarti, R. G. Laha and J.

Roy. Applied Statistics, 17(3):293–294, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1968:BRS

[353] O. L. Davies. Book reviews: Statistical Methods. 6th Edition, by George W. Snedecor and William G. Cochran. Applied Statistics, 17(3):294, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Disley:1968:BRI

[354] I. J. Disley. Book reviews: Industrial Production Models: A Theoretical Study, by Sven Dano. Applied Statistics, 17 (3):294–295, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1968:BRMb

[355] I. G. Evans. Book reviews: Multivariate Statistical Methods, by Donald F. Morrison. Applied Statistics, 17(3):295–296, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:1968:BRE

[356] P. L. Goldsmith. Book reviews: Elementary Sampling Theory, by Taro Yamane. Applied Statistics, 17(3):296, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:1968:BRP

[357] P. L. Goldsmith. Book reviews: Principles of Statistics. 2nd Edition, by M. G. Bulmer. Applied Statistics, 17(3): 296–297, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:1968:BRM

[358] P. L. Goldsmith. Book reviews: Modern Elementary Statistics. 3rd Edition, by John E. Freund. Applied Statistics, 17(3):297–298, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lovell:1968:BRI

[359] K. Lovell. Book reviews: Introduction to Statistical Analysis and Inference for Psychology and Education, by S. J. Armore. Applied Statistics, 17(3):298–299, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mellor:1968:BRI

[360] P. Mellor. Book reviews: Industrial Scheduling Abstracts (1950–1966), by Samuel Eilon and John R. King. Applied Statistics, 17(3):299–300, 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pridmore:1968:BRS

[361] W. A. Pridmore. Book reviews: Statistics for Scientists and Engineers, by R. Lowell Wine. Applied Statistics, 17 (3):300–301, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bunday:1968:BRS

[362] B. D. Bunday. Book reviews: Statistics and the Manager: The Use of Statistics and Probability in Managerial Decisions, by P. G. Moore. Applied Statistics, 17(3):301, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1968:BMPc

[363] Anonymous. Back matter. Applied Statistics, 17(3):[1], 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1969:CTQ

[364] A. F. Bissell. Cusum techniques for quality control. Applied Statistics, 18(1):1– 30, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Woodman:1969:RRS

[365] R. C. Woodman. Replacement rules for single and multi-component equipment. Applied Statistics, 18(1):31–40, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conolly:1969:COF

[366] B. W. Conolly and N. Hadidi. A comparison of the operational features of conventional queues with a self-regulating system. Applied Statistics, 18(1):41–53, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1969:MST

[367] J. C. Gower and G. J. S. Ross. Minimum spanning trees and single linkage cluster analysis. Applied Statistics, 18(1):54–64, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swan:1969:CML

[368] A. V. Swan. Computing maximumlikelihood estimates for parameters of the normal distribution from grouped and censored data. Applied Statistics, 18(1):65–69, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [427].

Jackson:1969:FGL

[369] O. A. Y. Jackson. Fitting a gamma or log-normal distribution to fibrediameter measurements on wool tops. Applied Statistics, 18(1):70–75, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1969:DCD

[370] J. N. R. Jeffers. Data capture: Data capture — A review of some current procedures. Applied Statistics, 18(1):76–78, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fraser:1969:DCU

[371] A. I. Fraser. Data capture: The use of automatic data-loggers in forest research. Applied Statistics, 18(1):78–81, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thompson:1969:DCS

[372] R. Thompson and J. T. Wood. Data capture: A semi-automatic directrecording balance. Applied Statistics, 18(1):82–85, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heyes:1969:DCU

[373] J. Heyes. Data capture: The use of trace assessors in applied statistics. Applied Statistics, 18(1):85–88, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thorne:1969:DCR

[374] C. Thorne. Data capture: Recording numerical data on port-a-punch cards. Applied Statistics, 18(1):88–89, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1969:DCD

[375] D. K. Lindley and D. H. Stewart. Data capture: Data processing in forestry with the aid of lector. Applied Statistics, 18(1):89–98, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stewart:1969:DCP

[376] Duncan H. Stewart. Data capture: Portable magnetic-tape encoders for computer data. Applied Statistics, 18 (1):98–100, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoyle:1969:MNO

[377] M. H. Hoyle. Miscellanea: A note on the optimal sampling fraction in sampling with a “non-sampling” cost. Applied Statistics, 18(1):101–102, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ross:1969:SAAa

[378] G. J. S. Ross. Statistical algorithms: Algorithm AS 13: Minimum spanning tree. Applied Statistics, 18(1):103– 104, March 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/13.

Ross:1969:SAAb

[379] G. J. S. Ross. Statistical algorithms: Algorithm AS 14: Printing the minimum spanning tree. Applied Statistics, 18 (1):105–106, March 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/14.

Ross:1969:SAAc

[380] G. J. S. Ross. Statistical algorithms: Algorithm AS 15: Single linkage cluster analysis. Applied Statistics, 18(1):106– 110, March 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/15.

Swan:1969:SAAa

[381] A. V. Swan. Statistical algorithms: Algorithm AS 16: Maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 18(1): 110–114, March 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/16. See remarks

[434, 430].

Swan:1969:SAAb

[382] A. V. Swan. Statistical algorithms: Algorithm AS 17: The reciprocal of Mills’s ratio. Applied Statistics, 18(1):115– 116, March 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/17.

Gower:1969:SAR

[383] J. C. Gower. Statistical algorithms:

Remark AS R1: A remark on Algorithm AS 1 sub-routine package. Applied Statistics, 18(1):116–117, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/1. See

Anonymous:1969:SAA

[384] Anonymous. Statistical algorithms: Algorithms AS 6 and AS 7: Corrigenda. Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [329, 330, 429, 431].

Cooper:1969:SAAa

[385] B. E. Cooper. Statistical algorithms: Algorithm AS 3: The integral of Student’s t-distribution: Corrigenda. Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/3.

Cooper:1969:SAAb

[386] B. E. Cooper. Statistical algorithms: Algorithm AS 4: An auxiliary function for distribution integrals: Corrigenda. Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/4. See cor-

rigendum [422].

Cooper:1969:SAAc

[387] B. E. Cooper. Statistical algorithms: Algorithm AS 5: The integral of the noncentral t-distribution: Corrigenda. Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/5.

Anonymous:1969:VIP

[388] Anonymous. Volume information. Applied Statistics, 18(1):[3], 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bithell:1969:SHA

[389] J. F. Bithell. The statistics of hospital admission systems. Applied Statistics, 18 (2):119–129, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Colquhoun:1969:CET

[390] D. Colquhoun. A comparison of estimators for a two-parameter hyperbola. Applied Statistics, 18(2):130–140, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Craig:1969:ETI

[391] J. Craig. On the elementary treatment of index numbers. Applied Statistics, 18 (2):141–152, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fienberg:1969:PGA

[392] Stephen E. Fienberg. Preliminary graphical analysis and quasiindependence for two-way contingency tables. Applied Statistics, 18(2):153–168, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sowden:1969:CBV

[393] R. R. Sowden and J. R. Ashford. Computation of the bi-variate normal integral. Applied Statistics, 18(2):169–180, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harvey:1969:MGA

[394] R. A. Harvey. Miscellanea: A graphical aid to visual fitting of straight lines. Applied Statistics, 18(2):181–185, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shahani:1969:MSG

[395] A. K. Shahani. Miscellanea: A simple graphical test of association for large samples. Applied Statistics, 18 (2):185–190, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cole:1969:MYA

[396] A. J. Cole and R. M. Campbell. Miscellanea: Yet another conversational mode program. Applied Statistics, 18 (2):190–191, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carpenter:1969:RFO

[397] L. C. Carpenter. Reviews: Fundamentals of Operations Research, by R. L. Ackoff and M. W. Sasieni. Applied Statistics, 18(2):192, 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Welford:1969:RSD

[398] B. P. Welford. Reviews: The Science of Decision Making, by A. Kaufmann and R. Audley. Applied Statistics, 18 (2):192–193, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1969:RHV

[399] G. B. Hill. Reviews: Health and Vital Statistics, by B. Benjamin. Applied Statistics, 18(2):193–194, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mack:1969:RTR

[400] C. Mack. Reviews: The Theory of Random Clumping, by S. A. Roach. Applied Statistics, 18(2):194, 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Machin:1969:REP

[401] David Machin. Reviews: The Elements of Probability and Sampling, by Frank A. Friday. Applied Statistics, 18 (2):194–195, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conway:1969:RAS

[402] Freda Conway. Reviews: Applied Statistics. Statistics for the Social Scientists: Volume 2, by K. A. Yeomans. Applied Statistics, 18(2):195–196, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1969:SAAa

[403] J. C. Gower. Statistical algorithms: Algorithm AS 18: Evaluation of marginal means. Applied Statistics, 18(2):197– 199, June 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/18.

Gower:1969:SAAb

[404] J. C. Gower. Statistical algorithms: Algorithm AS 19: Analysis of variance for a factorial table. Applied Statistics, 18 (2):199–202, June 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/19.

Nelder:1969:SAA

[405] J. A. Nelder. Statistical algorithms: Algorithm AS 20: The efficient formation of a triangular array with restricted storage for data. Applied Statistics, 18(2):

203–206, June 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/20.

Thayer:1969:SAA

[406] R. P. Thayer and R. F. Storer. Statistical algorithms: Algorithm AS 21: Scale selection for computer plots. Applied Statistics, 18(2):206–208, June 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/21. See corrigenda [424] and remark [428].

Day:1969:EAA

[407] D. J. Day and Jennifer E. Dunn. Estimating the audience for advertising on the outside of London buses. Applied Statistics, 18(3):209–220, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Roth:1969:ESM

[408] Philip M. Roth and Richard A. Stewart. Experimental studies with multiple responses. Applied Statistics, 18 (3):221–228, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hadidi:1969:IOC

[409] N. Hadidi and B. W. Conolly. On the improvement of the operational characteristics of single-server queues by the use of a queue-length-dependent service mechanism. Applied Statistics, 18 (3):229–240, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Box:1969:PET

[410] M. J. Box. Planning experiments to test the adequacy of non-linear models. Applied Statistics, 18(3):241–248, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chambers:1969:CSF

[411] John M. Chambers. A computer system for fitting models to data. Applied Statistics, 18(3):249–263, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Foster:1969:MCA

[412] F. G. Foster and T. M. F. Smith. Miscellanea: The computer as an aid in teaching statistics. Applied Statistics, 18 (3):264–270, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cochran:1969:MUC

[413] W. G. Cochran. Miscellanea: The use of covariance in observational studies. Applied Statistics, 18(3):270–275, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

James:1969:MTB

[414] William H. James. Miscellanea: Testing for birth-order effects in the presence of birth limitation or reproductive compensation. Applied Statistics, 18 (3):276–281, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gales:1969:BRF

[415] Kathleen Gales. Book reviews: Fieldwork Methods in General Use. First Report by Working Party on Interviewing Methods. Applied Statistics, 18(3):282, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kingman:1969:BRC

[416] J. F. C. Kingman. Book reviews: Convergence of Probability Measures, by P. Billingsley. Applied Statistics, 18(3):282, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Claringbold:1969:SAA

[417] Peter J. Claringbold. Statistical algorithms: Algorithm AS 22: The interaction algorithm. Applied Statistics, 18(3): 283–287, September 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/22.

Gower:1969:SAAc

[418] J. C. Gower. Statistical algorithms: Algorithm AS 23: Calculation of effects. Applied Statistics, 18(3):287–290, September 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/23.

Cunningham:1969:SAA

[419] S. W. Cunningham. Statistical algorithms: Algorithm AS 24: From normal integral to deviate. Applied Statistics, 18(3):290–293, September 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/24.

Lagoo:1969:SAA

[420] V. K. Lagoo. Statistical algorithms: Algorithm AS 25: Classification of means from analysis of variance. Applied Statistics, 18(3):294–298, September 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/25.

Hill:1969:SAR

[421] I. D. Hill. Statistical algorithms: Remark AS R2: A remark on Algorithm AS 2 “The Normal Integral”. Applied Statistics, 18(3):299–300, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/2. See

[325].

Cooper:1970:SAC

[422] B. E. Cooper. Statistical algorithms:

Corrigendum: Algorithm AS 4: An auxiliary function for distribution integrals. Applied Statistics, 19(2):204, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/4. See

[386].

Jennrich:1971:SAR

[423] R. I. Jennrich and P. I. Sampson. Statistical algorithms: Remark AS R3: A remark on Algorithm AS 10. Applied Statistics, 20(1):117–118, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/10. See

[348, 425].

Thayer:1971:SAC

[424] R. P. Thayer and R. F. Storer. Statistical algorithms: Corrigenda: Algorithm AS 21: Scale selection for computer plots. Applied Statistics, 20(1):118, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/21. See [406, 428].

Cooper:1971:SAR

[425] B. E. Cooper. Statistical algorithms: Remark AS R4: A remark on Algorithm AS 10: “The Use of Orthogonal Polynomials”. Applied Statistics, 20(2):216, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/10. See [348, 423].

Hitchin:1973:SAR

[426] D. Hitchin. Statistical algorithms: Remark AS R8: A remark on algorithms AS 4 and AS 5: An auxiliary function for distribution integrals, and the integral of the non-central t- distribution. Applied Statistics, 22(3):428, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [327, 328].

Swan:1974:CCM

[427] A. V. Swan. Corrigendum: Computing maximum-likelihood estimates for parameters of the normal distribution from grouped and censored data. Applied Statistics, 23(2):202, 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [368].

Rowlands:1974:SAR

[428] G. J. Rowlands and Rita M. Pocock. Statistical algorithms: Remark AS R10: On scale selections for scatter diagrams. Applied Statistics, 23(2):248–249, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [406, 424].

Farebrother:1974:SAR

[429] R. W. Farebrother and G. Berry. Statistical algorithms: Remark AS R12: A remark on Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 23(3):477, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/6. See [329,

384, 431].

Swan:1977:SAR

[430] A. V. Swan. Statistical algorithms: Remark AS R21: A remark on Algorithm AS 16 maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 26(1):122, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/16. See

[381, 434].

Barrett:1978:SAR

[431] J. C. Barrett and M. J. R. Healy. Statistical algorithms: Remark AS R27: A remark on Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 27(3): 379–380, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/6. See [429, 329, 384].

Freeman:1982:SAR

[432] P. R. Freeman. Statistical algorithms:

Remark AS R44: A remark on AS 6 and AS 7: Triangular decomposition of a symmetric matrix and inversion of a positive semi-definite symmetric matrix. Applied Statistics, 31(3):336–339, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [329, 330].

Chou:1985:SARb

[433] Youn-Min Chou. Statistical algorithms: Remark AS R56: A remark on Algorithm AS 5: The integral of the noncentral t-distribution. Applied Statistics, 34(1):102, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/5. See [328].

Hill:1987:SAR

[434] I. D. Hill. Statistical algorithms: Remark AS R70: A remark on Algorithm AS 16. maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 36 (1):119–120, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/16. See [381, 430].